### v2

Abstract
Quantum Process Tomography (QPT) methods aim at identifying, i.e. estimating, a quantum process. QPT is a major quantum information processing tool, since it especially allows one to experimentally characterize the actual behavior of quantum gates, that may be used as the building blocks of quantum computers. We here consider unitary, possibly dense (i.e. without sparsity constraints) processes, which corresponds to isolated systems. Moreover, we develop QPT methods that are applicable to a significant number of qubits and hence to a high state space dimension, which allows one to tackle more complex problems. Using the unitarity of the process allows us to develop methods that first achieve part of QPT by performing an eigenanalysis of the estimated density matrix of a process output. Building upon this idea, we first develop a class of complete algorithms that are single-stage, in the sense that they use only one eigendecomposition. We then extend them to multiple-stage algorithms (i.e. with several eigendecompositions), in order to address high-dimensional state spaces while being less limited by the estimation errors made when using an arbitrary given Quantum State Tomography (QST) algorithm as a building block of our overall methods. We first propose two-stage methods and we then extend them to dichotomic methods, whose number of stages increases with the considered state space dimension. The relevance of our methods is validated with simulations. Single-stage and two-stage methods efficiently apply up to 13 qubits on a standard PC (with 16 GB of RAM). Multi-stage methods yield an even higher accuracy (Normalized Mean Square Error down to $10^{-3}$).
Index Terms:
eigenanalysis of density matrix, high-dimensional quantum state space, quantum computing, quantum gate characterization, quantum machine learning (supervised / semi-supervised / unsupervised), quantum process tomography, quantum state tomography, unitary quantum gate
[…]
=-15pt
I Introduction
Quantum process tomography (QPT) may be considered as the quantum counterpart of classical system identification [1], that is, system parameter estimation. It was mainly [^1] See also [1] p. 398 for the other earliest references. introduced in 1997 in [2] and then extended (together with quantum process learning) e.g. in [1, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. QPT is a major quantum data processing tool, since it especially allows one to characterize the actual behavior of quantum gates, that may be used as the building blocks of quantum computers. This gate-based (or circuit-based) approach builds upon the unitary transforms performed by quantum gates (see Chapter 4 of [1]). It is currently one of the most prominent quantum computing paradigms and various commercial products based on this paradigm are available. This e.g. includes IBM’s Qiskit environment (see [28, 29, 30] and especially [31] for the use of reversible unitary gates), Amazon’s Braket tool (see [32] especially for reversible transformations), and Mathworks’s software package (see [33] e.g. for their statement “Quantum gates represent reversible operations that transform the quantum state according to unitary matrices”). Due to the importance of the above-mentioned unitary processes, various methods dedicated to such processes have been proposed, for performing unitary quantum process estimation/learning (see e.g. [3, 22, 13, 23, 24, 25, 26, 27]) and e.g. for the related topic of the inversion and transposition of unknown unitary operations [34].
QPT is a “‘data-driven” approach, and can thus be considered as a type of quantum machine learning (QML) problem (see [35] for a taxonomy of QML problems), because the input-output transform performed by the considered process is generally inferred from the values of input quantum states and/or from measurement results for output quantum states of that process. More precisely, usual QPT methods require one to fix a priori and hence know the above input “values” (i.e. states) of the process and to use information about the corresponding output “values” (i.e. states), that is derived from measurements performed at the output of the considered process. This corresponds to the supervised version of machine learning, also called the non-blind version of machine learning [36]. In contrast, the other extreme case is the one which requires the most limited knowledge about the process input whereas output measurements are still used, and this corresponds to the unsupervised, or blind, QPT (BQPT) configuration [36]. BQPT methods were introduced in [37] and then especially developed in [38, 39, 36]. In that case, the individual input state values are completely unknown, that is, neither fixed a priori nor estimated by means of measurements: in [37, 38], the input consists of unknown unentangled deterministic-coefficients pure states (DCPS); in [39, 36], it consists of random-coefficient pure states (RCPS) and the information available about these RCPS is restricted to some of their global, i.e. statistical, properties. DCPS are the usual form of pure states of quantum physics, whereas RCPS were introduced in [40] and then discussed in various related papers, especially including [41].
Between the above two extreme cases, a recently introduced approach (see [22] and references therein) was claimed to be semi-blind (i.e. semi-supervised), because it does not require the values of the input quantum states to be fixed a priori, but it estimates them, by also performing measurements at the input of the process for part of the available copies of these input states. This method thus allows one to use almost any values for the input states, whereas various supervised QPT methods are more constraining because they are only applicable to a specific, predefined, set of input values.
As detailed below, the different parts and variants of the QPT algorithms proposed in the present paper cover all above three types of learning. Some of them require their input states to be known, so that these states may be fixed a priori (supervised learning) or estimated from measurements (semi-supervised learning). Other parts and variants of our algorithms use input states whose values are not known, but that are required to meet a given condition (unsupervised learning). Besides, part of the proposed algorithms use “pure states”, where we employ the term “pure states” in the above-defined usual sense, that is, DCPS. In contrast, other proposed algorithms use statistical mixtures.
Within the above framework, we focus on an original case which combines the following two features. First, the considered process is constrained to be unitary, which corresponds to addressing a system that is isolated from its environment. Second, we are mainly interested in the case when the process has a high dimension, because it involves a significant number of qubits. This allows one to tackle more complex problems. We stress that we do not set any other conditions on the considered process. In particular, we allow (but do not request) the matrix that represents that process in the considered basis to be a so-called dense, i.e. non-sparse, matrix (we here use the standard terminology: a sparse matrix is a matrix that has a large percentage of zeros; see e.g. [42]). In contrast, some approaches from the literature for QPT and the related topic of Hamiltonian estimation request sparsity (see e.g. [43, 12]), but this was considered to be less relevant than unitarity in [22] and we explained above when the unitarity assumption is relevant.
Concerning the “mathematical principles” of the proposed QPT methods, in this paper we show that one of the key ingredients of our approach is the eigendecomposition of the (estimated) density matrix of a mixed output state of the considered process. Before moving to technical details and a variety of required extensions of this idea in the subsequent sections, this major idea can here be outlined as follows. The first output density matrix of the considered process is here denoted as $\rho_{2}$. As discussed in Section II, this density matrix reads
$\rho_{2}=U\hskip 2.84526pt\rho_{1}\hskip 2.84526ptU^{{\dagger}}$ (1)
where $U$ is the unitary matrix that represents the process to be estimated in the considered basis, $^{{\dagger}}$ is the transconjugate (i.e. Hermitian transpose) and $\rho_{1}$ is the density matrix of the first input mixed state that yields the output state $\rho_{2}$. The idea is then that, when the density matrix $\rho_{1}$ is selected to be diagonal in the considered and fixed basis (computational basis), Eq. (1) is an eigendecomposition of $\rho_{2}$. Therefore, in practice, starting from an estimate of $\rho_{2}$ derived from measurements, its eigendecomposition may be hoped to yield a result directly related to an estimate of $U$. This idea will then have to be refined further in this paper because: 1) eigendecomposition yields indeterminacies (defined below) and 2) fluctuations in output density matrix estimation limit the performance of this preliminary approach and then lead us to extend it in this paper.
The above preliminary description shows that the type of approaches proposed in this paper has some connections with works that were previously reported both in the classical and quantum domains. The connections with classical approaches are discussed in Appendix A, whereas those with the quantum domain may be defined as follows. A few recent papers dealing with quantum data also consider (1) in the framework of QPT. However, they only have the following limited relationships with our investigation. Ref. [44] aims at deriving the number (and values) of quantum states that should be applied to the input of a quantum process to be able to identify that process. However, that paper does not focus on the same input values as in the present paper and, as also noted in [22], it does not provide an explicit QPT algorithm. Ref. [3] provides such algorithms, that are different from ours and that are tested only for a Hilbert space dimension $d=5$, whereas we consider dimensions up to $d=2^{13}\simeq 8000$.
QPT algorithms may also be split in two classes, from the following point of view. Part of them first use the results of measurements at the output of the system by explicitly performing quantum state tomography (QST). This means that they estimate the ket (for a pure state) or the density matrix (for a statistical mixture) that represents that ouput state. They then use these estimated output states (and the corresponding input states, their estimates or their properties) so as to infer the considered process. Other QPT methods instead directly use the results of the considered types of measurements. In the present paper, we consider the first class of QPT methods and we focus on the part dedicated to QPT itself, as opposed to QST, as detailed in Section V.
The remainder of this paper is organized as follows. Section II is devoted to the introduction of several variants of our basic type of QPT methods, that is, single-stage methods. These methods have a limitation, that is defined in Section III. Therefore, we first introduce the two-stage extensions of the above methods in Section III, and then their general multi-stage form in Section IV. Numerical results illustrate the performance of the proposed methods in Section V and conclusions are drawn from this investigation in Section VI.
II Proposed single-stage QPT methods
As stated above, in all the configurations considered in this paper, the quantum system is assumed to be isolated from its environment (the proposed approaches may therefore also be used as an approximation for systems that have a limited interaction with their environment and that therefore lead to near-unitary processes, as in [3]). It is well-known that, in that case, the temporal evolution of the state of that system between given initial and final times is governed by a unitary operator (see [1, 45]; see also [46, 47] for the relationship with the system’s Hamiltonian). More precisely, using the above-defined notations, we consider a (first) mixed state defined by a density matrix $\rho_{1}$ at an initial time. As explained e.g. in [1, 45], the state of that system at a subsequent, final, time is then defined by a density matrix $\rho_{2}$ that may be expressed as in (1), where $U$ is a unitary matrix, that depends on the dynamics of the considered system.
The algorithms proposed in this section to estimate $U$ then consist of two parts. Each part involves a given type of input quantum state and an algorithm for exploiting the associated output quantum state (and possibly the input state). Each of the following subsections describes one of these two parts.
II-A First part of the methods
We here consider the situation when the first state applied to the analyzed process is a given statistical mixture, represented by the density matrix $\rho_{1}$. The process output is then also a mixed state and its density matrix $\rho_{2}$ may be expressed as (1). More specifically, we here introduce a first variant, in which the matrix $\rho_{1}$ is diagonal (see Appendix B about how to create it), whereas Appendix G addresses the more general case when this constraint is removed.
The matrix $\rho_{2}$ is known to be Hermitian, since it is a density matrix (see e.g. [45] p. 283). Therefore, it is normal (see [48] p. 100;  [^2] Throughout this paper, the reader may refer to [49], instead of [48].) and hence unitarily diagonalizable: see Theorem 1 and comments in its proof (all theorems cited in the core of this paper are provided in Appendix F). The first part of the QPT method proposed below is based on the eigendecomposition of $\rho_{2}$ (because (1) is such an eigendecomposition, as detailed below). More precisely, in practice, what is initially available for performing an eigendecomposition is not the theoretical matrix $\rho_{2}$ but an estimate $\widehat{\rho}_{2}$ of $\rho_{2}$, provided by a QST algorithm (we here use standard notations, e.g. as in [50, 51, 52]: an estimate of an unknown, deterministic, scalar, vector or matrix $\theta$ is denoted as $\widehat{\theta}$). Therefore, our QPT method preferably starts with two preprocessing steps, that consist of first deriving a Hermitian estimate $\widehat{\rho}_{3}$ from $\widehat{\rho}_{2}$ and then a unit-trace estimate $\widehat{\rho}_{4}$ from $\widehat{\rho}_{3}$ (see details in Appendix C), unless the considered QST algorithm already guarantees that $\widehat{\rho}_{2}$ meets these two properties.
Then, the core of our QPT algorithm performs an eigendecomposition of the Hermitian matrix $\widehat{\rho}_{4}$. This eigendecomposition problem has several solutions, corresponding to the indeterminacies of eigendecomposition detailed below. Any of these solutions is defined by two matrices, hereafter denoted as $D_{4}$ and $V_{4}$. $D_{4}$ is a diagonal matrix that contains the eigenvalues of $\widehat{\rho}_{4}$ in a arbitrary order (see e.g. [48] p. 102). Moreover, each column of $V_{4}$ is an eigenvector of $\widehat{\rho}_{4}$ (see (2) and Theorem 2) and $V_{4}$ is here unitary (because $\widehat{\rho}_{4}$ is unitarily diagonalizable), hence $V_{4}^{-1}=V_{4}^{{\dagger}}$. These results $V_{4}$ and $D_{4}$ of the eigendecomposition are linked to the initial matrix $\widehat{\rho}_{4}$ as follows:
$\widehat{\rho}_{4}$ $=$ $V_{4}\hskip 2.84526ptD_{4}\hskip 2.84526ptV_{4}^{-1}$ (2)
$=$ $V_{4}\hskip 2.84526ptD_{4}\hskip 2.84526ptV_{4}^{{\dagger}}.$ (3)
Comparing (3) to (1), with $\widehat{\rho}_{4}$ an estimate of $\rho_{2}$ and with $\rho_{1}$ a diagonal matrix, shows that (1) is one of the possible eigendecompositions of $\widehat{\rho}_{4}$ up to estimation errors. Therefore, the matrices $D_{4}$ and $V_{4}$ obtained in practice with our algorithm are respectively equal to $\rho_{1}$ and $U$ up to the indeterminacies of the eigendecomposition process (and up to estimation errors). In particular, $V_{4}$ is the first obtained quantity related to the estimation of $U$ (it could therefore be denoted as $\widehat{U}_{1}$). We now define its indeterminacies in more detail, and we show how to solve them.
In our first type of QPT methods described here, we constrain all diagonal values of the user-defined matrix $\rho_{1}$ to be different. This means that all eigenvalues in the eigendecompositions of $\rho_{2}$ and hence $\widehat{\rho}_{4}$ (ignoring estimation errors) are different. Therefore, each of them is associated with a different one-dimensional eigenspace. In other words, all eigenvectors corresponding to a given eigenvalue are along the same direction and the directions associated with two different eigenvalues are different (this is true for any complex square matrix: see [48] p. 47) and even orthogonal: see Theorem 3. Each eigenvalue is thus coupled with a single eigenvector in any considered eigendecomposition. The first indeterminacy of the eigendecomposition of $\widehat{\rho}_{4}$ is then a possible coupled permutation of 1) the columns of the matrix $V_{4}$ obtained as a first quantity towards the estimation of $U$ in this eigendecomposition and 2) the values along the diagonal of the matrix $D_{4}$, due to the above-mentioned arbitrary order of the latter diagonal values. This permutation problem may be solved as follows. As stated above, we constrain all diagonal values of $\rho_{1}$ to be different (at this stage of the discussion, we restrict ourselves to a theoretical approach, where we only need the eigenvalues to be different; further in this paper, we will discuss the gap required between them, when considering practical implementations of our approach: see Appendix D for the method considered in the present section). Moreover, we constrain all diagonal values of $\rho_{1}$ to be placed on that diagonal according to a given order. Hereafter, we will consider the case when they are in decreasing order. With respect to the above-defined machine learning terminology, this part of our algorithm is thus unsupervised: it does not require the exact value of the input state, i.e. the input density matrix $\rho_{1}$, to be known but it sets limited conditions on some of its properties.
The above known order of the diagonal values of $\rho_{1}$ is used to postprocess $D_{4}$, by permuting the values obtained on its diagonal so that they are in the same order as the order that we impose on $\rho_{1}$. More precisely, in the case considered hereafter, we permute the values obtained on the diagonal of $D_{4}$ so that they are in decreasing order (if some eigenvalues obtained in practice are nonreal, they are ordered according to their moduli). Since we now know which permutation is required to this end, we then use that specific permutation to reorder the columns of $V_{4}$ in the same way. The matrix with reordered columns thus obtained is denoted as $\widehat{U}_{2}$, since it is the second quantity that we build for estimating $U$.
Thanks to the above approach, the only indeterminacy that remains in $\widehat{U}_{2}$ with respect to $U$ is an unknown complex scale factor in each of its columns, since any such column belongs to the adequate one-dimensional eigenspace and is therefore defined up to a complex scale factor. Moreover, we create $\widehat{U}_{2}$ so that each of its columns has unit norm. To this end, if the considered eigendecomposition algorithm does not directly guarantee that this unit-norm condition is met, we postprocess the value of $\widehat{U}_{2}$ that it yields, by dividing each column of $\widehat{U}_{2}$ by its norm. After this postprocessing, the scale factor of each column of $\widehat{U}_{2}$, with respect to the column of $U$ that has the same index, reduces to a unit-modulus factor, i.e. a phase factor, because $U$ is unitary and therefore all its columns have unit norm, as those of $\widehat{U}_{2}$. $\widehat{U}_{2}$ then reads
$\widehat{U}_{2}=\left[e^{i\phi_{1}}u_{1},\dots,e^{i\phi_{d}}u_{d}\right]$ (4)
where $u_{1},\dots,u_{d}$ are the column vectors of the actual matrix $U$ and $\phi_{1},\dots,\phi_{d}$ are the phase indeterminacies with respect to the above vectors that remain at this stage. Denoting $D_{\phi}$ the diagonal matrix that contains the unknown phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$ on its main diagonal, (4) reads
$\widehat{U}_{2}=UD_{\phi}.$ (5)
The second part of our algorithm, presented below in Section II-B, aims at determining all phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$, up to a global phase factor (a process matrix $U$ can only be defined up to a single, i.e. global, phase factor).
II-B Second part of the methods
Thanks to the limited number of unknowns (namely the phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$) that remain in the matrix $\widehat{U}_{2}$ of (4) obtained in Section II-A, a single and pure state is then sufficient for estimating these unknowns. This yields the first variant of the second part of our methods, presented here. However, one may instead use a mixed state for estimating these unknowns, as shown in Appendix H.
We therefore here apply a known pure state, defined by a ket $\lvert\Psi_{1}\rangle$, at the input of the process (this part of the algorithm is therefore supervised, or semi-supervised if $\lvert\Psi_{1}\rangle$ is initially unknown but estimated by performing measurements for part of the available copies of that state; for the supervised case, Appendix E explains how $\lvert\Psi_{1}\rangle$ is chosen). The corresponding pure state at the output of that process is defined by the ket
$\lvert\Psi_{2}\rangle=U\lvert\Psi_{1}\rangle.$ (6)
What is available in practice is a ket $\lvert\widehat{\Psi}_{2}\rangle$, obtained by performing a QST at the output of the process, and equal to $\lvert\Psi_{2}\rangle$ up to an unknown phase factor and to estimation errors (see again Appendix C concerning a possible normalization of the QST output used here). When neglecting the above estimation errors,
$\lvert\widehat{\Psi}_{2}\rangle$ $=$ $e^{i\theta}\lvert\Psi_{2}\rangle$ (7)
$=$ $e^{i\theta}U\lvert\Psi_{1}\rangle.$ (8)
Our algorithm computes the ket
$\lvert\Psi_{3}\rangle=\widehat{U}_{2}^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle,$ (9)
based on the following motivation: combining (9), (5), (8) and considering that $U$ is unitary (hence $U^{{\dagger}}U=I$) yields
$\lvert\Psi_{3}\rangle=e^{i\theta}D_{\phi}^{*}\lvert\Psi_{1}\rangle$ (10)
where  stands for conjugation. This equation is very useful because, apart from the global phase $\theta$, its only unknown is $D_{\phi}$, i.e. the set of phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$ to be determined. This problem is easily solved by considering each component $k$, denoted as |.$\lvert.\rangle_{k}$, of the above kets. Using a ket $\lvert\Psi_{1}\rangle$ with nonzero components, (10) yields
$\frac{\lvert\Psi_{3}\rangle_{k}}{\lvert\Psi_{1}\rangle_{k}}=e^{i(\theta-\phi_{k})}\hskip 28.45274pt\forall k\in\{1,\dots,d\}.$ (11)
The associated matrix form reads
$\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)=e^{i\theta}D_{\phi}^{*}$ (12)
[…]
Based on the above properties, the final step of our algorithm creates, as follows, a new estimate of $U$ denoted as $\widehat{U}_{5}$ (the intermediate notations $\widehat{U}_{3}$ and $\widehat{U}_{4}$ are not used here but are required for the two-stage methods of Section III):
$\widehat{U}_{5}=\widehat{U}_{2}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle).$ (13)
This approach is used because, due to (12) and (5), Eq. (13) yields
$\widehat{U}_{5}=e^{i\theta}U$ (14)
which means that $\widehat{U}_{5}$ succeeds in restoring $U$, up to a global phase factor (and again up to the estimation errors that were mentioned above but implicit in the above equations).
A pseudo-code of the version of our algorithm corresponding to all this Section II is provided in Algorithm 1. This algorithm is called EQPT1, because it is our first proposed Eigenanalysis-based QPT algorithm.
Input : a) Estimate $\widehat{\rho}_{4}$ of output density matrix $\rho_{2}$ (provided by Quantum State Tomography (QST) and obtained for input density matrix (40)). b) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
Output : Estimate $\widehat{U}_{5}$ of quantum process matrix $U$.
begin
       /* Exploit $\widehat{\rho}_{4}$: */
       Eigendecomposition: derive 1) a diagonal matrix $D_{4}$ that contains the eigenvalues of $\widehat{\rho}_{4}$ in an arbitrary order and 2) a matrix $V_{4}$ whose columns are eigenvectors of $\widehat{\rho}_{4}$ in the same order as eigenvalues;
       Reorder the eigenvalues in $D_{4}$ in decreasing order and apply the same permutation to the columns of $V_{4}$ to create the matrix $\widehat{U}_{2}$;
       /* If the above eigendecomposition algorithm does not yield unit-norm eigenvectors, then create them as follows: */
       Divide each column of $\widehat{U}_{2}$ by its norm;
       /* Exploit $\lvert\widehat{\Psi}_{2}\rangle$: */
       $\lvert\Psi_{3}\rangle=\widehat{U}_{2}^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}=\widehat{U}_{2}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
end
Algorithm 1 EQPT1: first Eigenanalysis-based Quantum Process Tomography algorithm (composed of one stage). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see Appendix C.
III Proposed two-stage QPT methods
III-A Motivation
The dimension $d$ of the state space (and hence the dimension $d\times d$ of the considered process matrix $U$) that can be addressed by the type of QPT methods proposed in Section II is limited by the following phenomenon. As explained in that section, these methods require all eigenvalues of the input density matrix $\rho_{1}$ of the process to be different. Moreover, they should be separated by a minimum gap to be distinguishable, so that their estimates are not permuted in practice (in order to reorder them and the associated eigenvectors correctly). Besides, all of them range from 0 to 1, because they are nonnegative and their sum is equal to one, as stated above. All these conditions yield an upper bound, denoted as $n_{max}$, for the acceptable number of eigenvalues of $\rho_{1}$. Since the number of (different) eigenvalues of $\rho_{1}$ is always equal to $d$ for the methods of Section II, this yields an upper bound for the state space dimension $d$ that can be handled by these methods, and this bound on $d$ is also equal to $n_{max}$:
$d\leq n_{max}.$ (15)
In the present section, we aim at addressing higher values of $d$. To this end, we move to the case when some eigenvalues of the input density matrix $\rho_{1}$ are equal and we therefore introduce extensions of the QPT methods of Section II to be able to handle that more complex case.
III-B First part of the methods
As in Section II-A, the method proposed here starts by considering a given mixed state that has a diagonal density matrix $\rho_{1}$ and that is applied to the input of the considered unitary process defined by a matrix $U$. However, we here accept to have identical values on the main diagonal of $\rho_{1}$, at the expense of then having to handle more complex eigendecomposition properties for the process output, as shown below. More precisely, each value that appears on the main diagonal of $\rho_{1}$ appears $d_{1}$ times, on adjacent rows, and that diagonal contains $d_{2}$ different values. The number of rows of $\rho_{1}$, equal to the dimension of the considered state space, is thus equal to
$d=d_{1}d_{2}$ (16)
(in practice, we aim at considering the case when $d_{1}$, $d_{2}$, and hence $d$ are powers of 2, with an application to qubits). If requesting the same minimum gap between different eigenvalues as in Section II-A, or in other words, if replacing the bound (15) on $d$ in Section II by the same bound value for $d_{2}$ instead of $d$ here, then the maximum value of the state space dimension (16) that can be handled here is $d_{1}$ times higher than the maximum dimension that can be addressed by the methods of Section II, which is our goal, ideally. However, it is not guaranteed that the methods proposed hereafter can handle the same gap between different eigenvalues as in Section II-A with the same accuracy, especially due to the more complex structure of these extended methods. The practical accuracy of all proposed methods is therefore investigated numerically further in this paper.
Denoting $r_{1},\dots,r_{d_{2}}$ the different eigenvalues of $\rho_{1}$, that chosen density matrix reads
$\rho_{1}=\left[\begin{tabular}[]{llllllll}$r_{1}$&0&$\dots$&&&&&0\\ &$\ddots$\\ &&$r_{1}$\\ &&&$r_{2}$\\ &&&&$\ddots$\\ &&&&&$r_{2}$\\ &&&&&&$\ddots$\\ 0&$\dots$&&&&&0&$r_{d_{2}}$\end{tabular}\right]$ (17)
with each diagonal value repeated $d_{1}$ times. This may also be expressed as
$\rho_{1}=\mathrm{diag}(r)\otimes I_{d_{1}}$ (18)
where $r$ is the vector that contains the values $r_{1},\dots,r_{d_{2}}$, whereas $I_{d_{1}}$ is the $d_{1}\times d_{1}$ identity matrix and $\otimes$ stands for the Kronecker product [53]. Without loss of generality, we here again request the values on the diagonal of $\rho_{1}$ to be in nonincreasing order: $r_{1}>r_{2}>\dots>r_{d_{2}}$. Moreover, we preferably use values that are uniformly distributed (that is, multiples of the same step, with zero excluded), as in the method of Section II (see details for the latter method in Appendix D). In the configuration of the present section, (40) is thus replaced by:
$r_{k}=\frac{2(d_{2}-k+1)}{d(d_{2}+1)}\hskip 14.22636ptk\in\{1,\dots,d_{2}\}.$ (19)
Here again, the process output corresponding to $\rho_{1}$ is a mixed state and its density matrix $\rho_{2}$ may be expressed as (1). Moreover, as in Section II-A, $\rho_{2}$ is Hermitian and hence unitarily diagonalizable, and it has unit trace. Besides, in practice, an estimate $\widehat{\rho}_{2}$ of $\rho_{2}$ is derived by a QST algorithm and, if required by the considered QST algorithm, we preprocess this estimate of $\rho_{2}$ in order to derive its Hermitian and unit-trace version, denoted as $\widehat{\rho}_{4}$. We then perform an eigendecomposition of $\widehat{\rho}_{4}$. The analysis of the properties of this eigendecomposition here requires additional care, because the considered matrix has various identical eigenvalues.
The properties of Section II-A that remain true here are as follows. The eigendecomposition is not unique, Eq. (2)-(3) still apply, $V_{4}$ and $D_{4}$ are respectively a unitary matrix and a diagonal matrix, the $k$th column of $V_{4}$ is an eigenvector of $\widehat{\rho}_{4}$ and the associated eigenvalue is the $k$th value on the diagonal of $D_{4}$ (see again Theorem 2 for the latter property).
In contrast, what is different here as compared with Section II-A is that the same eigenvalue here appears $d_{1}$ time and is thus associated with $d_{1}$ eigenvectors, instead of only one in Section II-A. Eigendecomposition properties are more complex than above in this general framework but, fortunately, their complexity is here limited by the fact that the matrix $\widehat{\rho}_{4}$ to be decomposed is Hermitian. For this class of matrices (and for all normal matrices), the so-called geometric multiplicity of each eigenvalue is guaranteed to be equal to its so-called algebraic multiplicity, i.e. these matrices are nondefective (see details in [48] p. 57, 58, 103). More explicitly, this means that each eigenvalue of $\widehat{\rho}_{4}$ is here known to appear $d_{1}$ times at arbitrary locations on the diagonal of $D_{4}$ (algebraic multiplicity), since these eigenvalues are the same, in a different order, as in another eigendecomposition of the same matrix $\widehat{\rho}_{4}$ (neglecting estimation errors), namely the eigendecomposition (1). Therefore, the dimension of the eigenspace associated with any eigenvalue of $\widehat{\rho}_{4}$ is also equal to $d_{1}$. So, when performing the eigendecomposition of $\widehat{\rho}_{4}$, the considered decomposition algorithm should provide $d_{1}$ linearly independent eigenvectors for each eigenvalue, and we hereafter assume that practical eigendecomposition algorithms succeed in finding them (that was indeed the case in all the tests reported further in this paper). As an overall result, when performing the eigendecomposition of $\widehat{\rho}_{4}$, 1) we get a diagonal matrix $D_{4}$ with $d=d_{1}d_{2}$ eigenvalues that may be in an arbitrary order, and with each encountered value repeated $d_{1}$ times and 2) we get a unitary matrix $V_{4}$ whose columns are eigenvectors that have the above-defined properties.
As in Section II-A, we then apply a joint permutation to the diagonal elements of $D_{4}$ and to the columns of $V_{4}$, so that the values corresponding to the diagonal elements of $D_{4}$ become in nonincreasing order. Considering the resulting complete set of values from the top to the bottom of the diagonal of the reordered version of $D_{4}$, we thus get $d_{2}$ successive subsets, with the same value repeated $d_{1}$ times in each subset (in practice, these $d_{1}$ values may be slightly different, due to estimation errors), and with different and decreasing values from one subset to the next one. In other words, up to estimation errors, the reordered version of $D_{4}$ is nothing but the matrix $\rho_{1}$ of (17). Each subset of identical values along the diagonal of the obtained matrix (values equal to eigenvalue $r_{m_{1}}$) has an index $m_{1}$, with $m_{1}\in\{1,\dots,d_{2}\}$, and consists of elements with indices [^3] In this whole paper except Appendix K, the index of the first row and column of a matrix is equal to 1.
$(m_{1}-1)d_{1}+n_{1},\hskip 14.22636pt\mathrm{with}\hskip 14.22636ptn_{1}\in\{1,\dots,d_{1}\}.$ (20)
Moreover, the corresponding matrix, derived from $V_{4}$, with reordered eigenvectors in its columns is here again denoted as $\widehat{U}_{2}$. Its subset of columns that also have the indices defined by (20) is a set of $d_{1}$ eigenvectors that form a basis of the $d_{1}$-dimensional eigenspace of $\widehat{\rho}_{4}$ associated with its eigenvalue $r_{m_{1}}$. This eigenspace, obtained in the first part of this method, is hereafter denoted as ${\cal S}_{1,m_{1}}$.
Importantly, ${\cal S}_{1,m_{1}}$ also has the following relationship with the actual matrix $U$ to be identified. Eigenvectors corresponding to distinct eigenvalues are orthogonal for a normal matrix (see Theorem 3) and hence especially for a Hermitian matrix (see also [48] p. 171). This has the following consequence, considering that (1) and (17) define an eigendecomposition of $\rho_{2}$: each eigenvalue $r_{m_{1}}$ of $\rho_{2}$, with a multiplicity $d_{1}$, corresponds to a $d_{1}$-dimensional eigenspace 1) that is orthogonal to the eigenspaces defined by all other eigenvalues and 2) that is the subspace defined by the columns of $U$ that have the indices defined by (20). But this eigenspace is nothing but ${\cal S}_{1,m_{1}}$, since $\rho_{2}$ is equal to $\widehat{\rho}_{4}$ (up to estimation errors). This shows that this first part of our second type of methods does not yet succeed in separately identifying each column of $U$, but it succeeds in finding a basis for each of the $d_{2}$ subspaces ${\cal S}_{1,m_{1}}$ that corresponds to a value $m_{1}$, with $m_{1}\in\{1,\dots,d_{2}\}$, and that is associated with the $d_{1}$ columns of $U$ that have the indices defined by (20). We hereafter show how to extend that method so as to identify each column of $U$ separately.
III-C Second part of the methods
The two important properties that we showed in Section III-B for the considered (possibly high) dimension $d=d_{1}d_{2}$ are as follows. First, we proved that we can “partition” the set of $d$ columns of $U$ in $d_{2}$ subsets, in the sense that, in Section III-B, we identified each $d_{1}$-dimensional subspace ${\cal S}_{1,m_{1}}$ spanned by the columns of $U$ that compose the subset with index $m_{1}$. Second, we proved that we can select which partition we perform: the columns of $U$ that are gathered in the same subset are those that correspond to the same eigenvalue of $\rho_{2}$ and hence $\widehat{\rho}_{4}$.
In the present section, we take advantage of these results to create another partition of the columns of $U$, in order to then jointly exploit these two partitions as explained further in Section III-D. We thus create a two-stage algorithm in the sense that it successively uses two eigendecompositions of density matrices estimated at the output of the process $U$, namely the eigendecompositions respectively described in Section III-B and in the present subsection. This should be contrasted with the algorithms of Section II, that are single-stage methods in the sense that they perform a single eigendecomposition.
More precisely, the input density matrix $\rho_{5}$ here applied to the considered unitary process is different from $\rho_{1}$ but, here again, it is diagonal, it contains $d_{2}$ different values that each appear $d_{1}$ times and these values are the same as in Section III-B. However, they are here placed in a different order: the set of values on the main diagonal of $\rho_{5}$ consists of a series of $d_{1}$ identical subsets, each of which successively contains the $d_{2}$ values $r_{1},\dots,r_{d_{2}}$. The density matrix $\rho_{5}$ thus reads
$\rho_{5}=\left[\begin{tabular}[]{lllllll}$r_{1}$&0&$\dots$&&&&0\\ &$\ddots$\\ &&$r_{d_{2}}$\\ &&&$r_{1}$\\ &&&&$\ddots$\\ &&&&&$r_{d_{2}}$\\ 0&&&&&&$\ddots$\end{tabular}\right].$ (21)
This may also be expressed as
$\rho_{5}=I_{d_{1}}\otimes\mathrm{diag}(r)$ (22)
with the same notations as in (18).
This input density matrix $\rho_{5}$ is used in the same way as in Section III-B. So, the density matrix $\rho_{6}$ of the corresponding process output may be derived from $\rho_{5}$ in the same way as in (1). An estimate $\widehat{\rho}_{6}$ of $\rho_{6}$ is derived by a QST algorithm and, if required by the considered QST algorithm, we preprocess $\widehat{\rho}_{6}$ in order to derive: 1) its Hermitian version $\widehat{\rho}_{7}$ (in the same way as in (38)) and then 2) its Hermitian and unit-trace version $\widehat{\rho}_{8}$ (in the same way as in (39)). We then perform an eigendecomposition of $\widehat{\rho}_{8}$. This yields a diagonal matrix $D_{8}$ containing its eigenvalues and a matrix $V_{8}$ whose columns are its associated eigenvectors.
We then apply a joint permutation to the diagonal elements of $D_{8}$ and to the columns of $V_{8}$, so that the values corresponding to the diagonal elements of $D_{8}$ become in nonincreasing order. Here again, each of the $d_{2}$ successive subsets of $d_{1}$ identical values along the diagonal of the reordered version of $D_{8}$ (values equal to eigenvalue $r_{m_{2}}$) has an index $m_{2}$, with $m_{2}\in\{1,\dots,d_{2}\}$, and consists of elements with indices
$(m_{2}-1)d_{1}+n_{2},\hskip 14.22636pt\mathrm{with}\hskip 14.22636ptn_{2}\in\{1,\dots,d_{1}\}.$ (23)
Moreover, the corresponding matrix, derived from $V_{8}$, with reordered eigenvectors in its columns is denoted as $\widehat{U}_{3}$  [^4] In this whole paper, all the quantities of importance related to estimates of the considered process are denoted with the same letter and increasing indices 1, 2 and so on. When describing our first type of methods in Section II, for the sake of clarity we used detailed notations for these estimates, therefore with a first notation (with index 1) for the quantity before column reordering and a second notation (with index 2) for the quantity after column reordering. However, we then did not use the first notation. We here want to make it clear that, for the second type of methods described in the present section, we also consider two such quantities, namely before and after reordering. However, we do not introduce a notation for the quantity before reordering because, here again, we would not use it afterwards. Therefore, we here only introduce a single quantity, with an index equal to 3.. Its subset of columns that also have the indices defined by (23) is a set of $d_{1}$ eigenvectors that form a basis of the $d_{1}$-dimensional eigenspace of $\widehat{\rho}_{8}$ associated with its eigenvalue $r_{m_{2}}$. This eigenspace, obtained in the second part of this method, is hereafter denoted as ${\cal S}_{2,m_{2}}$.
${\cal S}_{2,m_{2}}$ also has a relationship with the actual matrix $U$ to be identified, but it is different from the relationship disclosed for ${\cal S}_{1,m_{1}}$ in Section III-B. Here again, ${\cal S}_{2,m_{2}}$ is the $d_{1}$-dimensional subspace spanned by the $d_{1}$ columns of $U$ that correspond to the eigenvalue $r_{m_{2}}$. However, here the indices of these columns are different from those of Section III-B: they are equal to the column indices of all occurrences of $r_{m_{2}}$ in (21), that is
$m_{2}+n_{2}d_{2}\hskip 14.22636pt\mathrm{with}\hskip 14.22636ptn_{2}\in\{0\dots,d_{1}-1\}.$ (24)
This shows that this second part of our second type of methods succeeds in finding a basis for each of the $d_{2}$ subspaces ${\cal S}_{2,m_{2}}$ that: 1) corresponds to a value $m_{2}$, with $m_{2}\in\{1,\dots,d_{2}\}$, and 2) is associated with the $d_{1}$ columns of $U$ that have the indices defined by (24). We hereafter show how to combine the properties of these subspaces ${\cal S}_{2,m_{2}}$ and of the subspaces ${\cal S}_{1,m_{1}}$ derived in the first part of this algorithm.
III-D Third part of the methods
We now consider all subspace intersections ${\cal S}_{1,m_{1}}\cap{\cal S}_{2,m_{2}}$, for $m_{1}\in\{1,\dots,d_{2}\}$ and $m_{2}\in\{1,\dots,d_{2}\}$. The subspaces ${\cal S}_{1,m_{1}}$ and ${\cal S}_{2,m_{2}}$ are each defined by a set of column vectors of $U$ and we want each such pair of subspaces to share only one column vector (or possibly no columns for some of these intersections). When this condition is met, ${\cal S}_{1,m_{1}}\cap{\cal S}_{2,m_{2}}$ is equal to the one-dimensional subspace defined by this shared column vector, because $U$ is unitary and all its column vectors are therefore orthogonal (see [48] p. 67). Moreover, we want the set of $d_{2}^{2}$ intersection column vectors thus created to yield all $d$ columns of $U$ (up to scale factors). This requires $d_{2}^{2}\geq d$ and hence, due to (16):
$d_{1}\leq d_{2}.$ (25)
Besides, the number $d_{2}$ of different eigenvalues of the input density matrix (as defined in Section III-B) is upper bounded by the value $n_{max}$, as explained above. Therefore, due to (25) and (16), the state space dimension $d$ that can be handled by the methods proposed in the present section is upper bounded by:
$d\leq(n_{max})^{2}.$ (26)
This maximum space dimension is therefore much higher than the bound (15) faced with the methods of Section II, under the assumption that $n_{max}$ here remains the same as with single-stage methods, as discussed above. In that case, the methods of the present section are much more attractive for handling high-dimensional spaces (but more complex).
In the remainder of this section, we only consider the case when
$d_{1}=d_{2}$ (27)
which makes the analysis simpler. Combining this condition with (16) yields
$d=d_{1}^{2}=d_{2}^{2}.$ (28)
When $d_{2}$ is moreover set to its maximum value $n_{max}$, we obtain $d=(n_{max})^{2}$, so that we reach the upper bound defined by (26).
One then analyzes how the indices of the columns of $U$ corresponding to ${\cal S}_{1,m_{1}}$ (defined by (20)) are interleaved with the indices of the columns of $U$ corresponding to ${\cal S}_{2,m_{2}}$ (defined by (24)), for any given couple of values $(m_{1},m_{2})$, with $m_{1}\in\{1,\dots,d_{2}\}$ and $m_{2}\in\{1,\dots,d_{2}\}$. This shows that, whatever $(m_{1},m_{2})$, the sets of columns of $U$ respectively associated with ${\cal S}_{1,m_{1}}$ and ${\cal S}_{2,m_{2}}$ share one and only one column, and that its index is
$(m_{1}-1)d_{1}+m_{2}.$ (29)
Moreover, when $m_{1}$ and $m_{2}$ are varied over all their ranges $m_{1}\in\{1,\dots,d_{2}\}$ and $m_{2}\in\{1,\dots,d_{2}\}$, the resulting set of $d_{2}^{2}$ indices (29) spans all $d=d_{1}d_{2}=d_{2}^{2}$ columns of $U$. This approach therefore succeeds in identifying all columns of $U$, up to scale factors at this stage.
The practical algorithm proposed for implementing the above approach operates as follows. Successively for each couple $(m_{1},m_{2})$, with $m_{1}\in\{1,\dots,d_{2}\}$ and $m_{2}\in\{1,\dots,d_{2}\}$, we need to determine a vector situated in the 1-dimensional intersection of the subspaces ${\cal S}_{1,m_{1}}$ and ${\cal S}_{2,m_{2}}$. The data available in practice for defining ${\cal S}_{1,m_{1}}$ consist of the set of columns of $\widehat{U}_{2}$ that have the indices defined by (20) and the data that define ${\cal S}_{2,m_{2}}$ consist of the set of columns of $\widehat{U}_{3}$ that have the indices defined by (23). To determine this 1-dimensional subspace intersection, we here use the first direction provided by the standard Canonical Correlation Analysis, or CCA, e.g. described in [54] and directly available in the Matlab scientific software [55]. This approach is by the way closely related to the eigendecomposition tool, that is the core of the methods proposed in the present paper [^5] It should also be noted that an alternative to CCA, based on Singular Value Decomposition, and that allows one to determine subspace intersections, is available in [56]. It is a building block of the method called SIBIS, for “Subspace-Intersection Blind Identification and Separation”. Beyond its subspace-intersection tool, SIBIS has the following relationships with the present paper in terms of final goal. SIBIS performs 1) Blind Mixture Identification, or BMI [57, 58, 59, 60], which is a type of classical system identification problem and is thus related to (the classical counterpart of) QPT and 2) Blind Source Separation, which is a problem closely connected to classical BMI..
By calculating the intersection ${\cal S}_{1,m_{1}}\cap{\cal S}_{2,m_{2}}$, we obtain one column vector of $U$, possibly up to an arbitrary complex factor, unless this vector is guaranteed to be normalized by the considered CCA tool. If it is not, we then reduce this factor to a phase factor, by dividing this intersection vector by its norm, so as to enforce its norm to become equal to one (again because $U$ is known to be unitary and thus to have unit-norm column vectors).
All these normalized results of CCA are used to create an additional matrix related to the estimation of $U$ and therefore denoted as $\widehat{U}_{4}$: each application of CCA to a given couple $(m_{1},m_{2})$ yields a normalized vector that is stored in the column of $\widehat{U}_{4}$ whose index is defined by (29). The matrix $\widehat{U}_{4}$ obtained at this stage is therefore equal to $U$ up to an arbitrary phase factor in each of its columns (and up to estimation errors).
III-E Fourth part of the methods
The properties of $\widehat{U}_{4}$ defined at the end of the previous section are exactly the same as those obtained in our single-stage methods for $\widehat{U}_{2}$: see Section II-A, including (4). Therefore, the final stage of the method considered here consists of processing $\widehat{U}_{4}$ in the same way as we processed $\widehat{U}_{2}$ in Section II-B. This yields a matrix $\widehat{U}_{5}$ that here again succeeds in restoring $U$, up to only a global phase factor.
A pseudo-code of the version of our algorithm described in this Section III is provided in Algorithm 2. This algorithm is called EQPT2. Two variants (called EQPT3 and EQPT4) of two-stage QPT methods are moreover presented in Appendix J.
Input : a) Estimate $\widehat{\rho}_{4}$ of output density matrix $\rho_{2}$ (provided by QST and obtained for input density matrix (17) with the diagonal values defined in (19)). b) Estimate $\widehat{\rho}_{8}$ of output density matrix $\rho_{6}$ (provided by QST and obtained for input density matrix (21) with the diagonal values defined in (19)). c) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
Output : Estimate $\widehat{U}_{5}$ of quantum process matrix $U$.
begin
       /* Exploit $\widehat{\rho}_{4}$: */
       Eigendecomposition: derive 1) a diagonal matrix $D_{4}$ that contains the eigenvalues of $\widehat{\rho}_{4}$ in an arbitrary order and 2) a matrix $V_{4}$ whose columns are eigenvectors of $\widehat{\rho}_{4}$ in the same order as eigenvalues;
       Reorder the eigenvalues in $D_{4}$ in nonincreasing order and apply the same permutation to the columns of $V_{4}$ to create the matrix $\widehat{U}_{2}$;
       /* Exploit $\widehat{\rho}_{8}$: */
       Eigendecomposition: derive 1) a diagonal matrix $D_{8}$ that contains the eigenvalues of $\widehat{\rho}_{8}$ in an arbitrary order and 2) a matrix $V_{8}$ whose columns are eigenvectors of $\widehat{\rho}_{8}$ in the same order as eigenvalues;
       Reorder the eigenvalues in $D_{8}$ in nonincreasing order and apply the same permutation to the columns of $V_{8}$ to create the matrix $\widehat{U}_{3}$;
       /* Combine $\widehat{U}_{2}$ and $\widehat{U}_{3}$: */
       for $m_{1}=1$ to $d_{2}$ do
             for $m_{2}=1$ to $d_{2}$ do
                   Set column of $\widehat{U}_{4}$ with index equal to $((m_{1}-1)d_{1}+m_{2})$: Set it to first vector provided by CCA of 1) the set of columns of $\widehat{U}_{2}$ with indices $((m_{1}-1)d_{1}+n_{1})$ with $n_{1}\in\{1,\dots,d_{1}\}$ and 2) the set of columns of $\widehat{U}_{3}$ with indices $((m_{2}-1)d_{1}+n_{2})$ with $n_{2}\in\{1,\dots,d_{1}\}$;
                   /* If the CCA tool does not provide unit-norm vectors, then do it: */
                   Divide above column of $\widehat{U}_{4}$ by its norm;
             end for
       end for
      /* Exploit $\lvert\widehat{\Psi}_{2}\rangle$: */
       $\lvert\Psi_{3}\rangle=\widehat{U}_{4}^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}=\widehat{U}_{4}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
end
Algorithm 2 EQPT2: second Eigenanalysis-based QPT algorithm (composed of two stages). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see text.
IV Proposed multi-stage QPT methods
The QPT methods proposed in the previous sections can be extended as follows, by selecting density matrices $\rho_{1}$ of input mixed states that are diagonal, that contain several occurrences of each value that appears on their diagonal and that have additional properties. More precisely, here again, each value that appears on the main diagonal of $\rho_{1}$ appears $d_{1}$ times, with $d_{1}>1$, and that diagonal contains $d_{2}$ different values, here possibly with $d_{2}\neq d_{1}$. The number of rows of $\rho_{1}$, equal to the dimension of the considered state space, is thus here again defined by (16). The new feature introduced here consists of storing all $d_{1}$ occurrences of any given diagonal value of $\rho_{1}$ in $b_{n}$ blocks of adjacent rows, with each block consisting of $b_{s}$ rows, and hence with
$b_{n}b_{s}=d_{1}.$ (30)
Besides, we consider the case when $b_{n}$, $b_{s}$ and $d_{1}$ are powers of 2. We denote as ${b_{\ell}}$ the integer-valued logarithm of $b_{n}$ in base 2: we have
$b_{n}=2^{b_{\ell}}.$ (31)
We hereafter consider all possible integer values of ${b_{\ell}}$, for a fixed value $d_{1}$ and varying values of $b_{n}$ and $b_{s}$. In other words, we consider all linked powers of 2 for $b_{n}$ and $b_{s}$, in their acceptable range. The minimum of ${b_{\ell}}$ is 0, corresponding to the minimum of $b_{n}$, equal to 1, and to the maximum of $b_{s}$, equal to $d_{1}$. Similarly, the maximum of ${b_{\ell}}$ is equal to $\log_{2}d_{1}$, corresponding to the maximum of $b_{n}$, that is equal to $d_{1}$, while $b_{s}$ reaches its minimum, equal to 1.
Moreover, these blocks are organized as follows: from the top to the bottom of the diagonal of $\rho_{1}$, we have a block of values equal to $r_{1}$, then a block of values $r_{2}$, and so on until a block of values $r_{d_{2}}$, and then the same structure again and again, that is, blocks of values $r_{1},r_{2},\dots,r_{d_{2}}$ (when there exist several blocks containing the same value, i.e. when $b_{n}>1$). Here again, we request these possible values to be defined in decreasing order, i.e. $r_{1}>r_{2}>\dots>r_{d_{2}}$, and to take the same values as in (19). The two extreme cases of this general framework are nothing but the two specific cases that we already used for the above-defined two-stage methods: when setting ${b_{\ell}}=0$ and hence $b_{n}=1$ and $b_{s}=d_{1}$, we obtain the density matrix of (17) whereas with ${b_{\ell}}=\log_{2}d_{1}$ and hence $b_{n}=d_{1}$ and $b_{s}=1$, we get the density matrix of (21).
The main idea of the above-defined two-stage methods was to exploit intersections of two subspaces of eigendecompositions resulting from the above density matrices (17) and (21). Here, we extend this idea to general multi-stage methods, with $(\log_{2}d_{1}+1)$ stages. To this end, we consider all $(\log_{2}d_{1}+1)$ input density matrices that belong to the above-defined general framework, by considering all $(\log_{2}d_{1}+1)$ possible values of ${b_{\ell}}$, i.e. all integers from 0 to $\log_{2}d_{1}$. Using the same approach as in Section III (that is, performing the eigendecomposition and reordering for the estimated output density matrix), for each value of ${b_{\ell}}$ and hence of $b_{n}$ and $b_{s}$, we obtain a set of $d_{2}$ subspaces denoted as ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$, with $m_{{b_{\ell}}}\in\{1,\dots,d_{2}\}$ (each index $m_{{b_{\ell}}}$ is used to define each subspace in the eigendecomposition of the stage ${b_{\ell}}$ of the algorithm proposed here; its correspondence with the structure of the input density matrix is not the same as when we used the indices $m_{1}$ and $m_{2}$ in Section III). These subspaces ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$ are defined in more detail in Appendix K. We then extend the approach of Section III by performing subspace intersections as follows. We successively consider all $d_{2}^{(\log_{2}d_{1}+1)}$ possible values of the set of $(\log_{2}d_{1}+1)$ subspace indices $m_{{b_{\ell}}}$ (with $m_{{b_{\ell}}}\in\{1,\dots,d_{2}\}$), i.e. we consider all possible values of the set of indices $\{m_{0},\dots,m_{\log_{2}d_{1}}\}$. For each such set of indices, we derive the intersection of the corresponding $(\log_{2}d_{1}+1)$ subspaces ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$. We thus form $d_{2}^{(\log_{2}d_{1}+1)}$ subspace intersections. As in Section III-D, we want each of these intersections to be one-dimensional (or possibly to be restricted to the null vector for some of these intersections). In order to identify all $d$ columns of $U$, the number of subspace intersections should be at least equal to $d$. Using (16), this condition reads
$d_{2}^{\log_{2}d_{1}}d_{2}\geq d_{1}d_{2}.$ (32)
Taking the logarithm, in base 2, of that equation yields
$\log_{2}d_{2}\geq 1$ (33)
that is
$d_{2}\geq 2.$ (34)
In particular, for $d_{2}=2$, we have equalities in (34) back to (32) and the number $d_{2}^{(\log_{2}d_{1}+1)}$ of intersections is exactly equal to the dimension (16). It should be noted that, unlike in the two-stage methods, we here do not face the constraints (25) and hence (26). This is because we pay the price by using a multi-stage, hence more complex, approach: the method proposed here potentially applies to any state space dimension $d$ and any number $d_{2}$ of different eigenvalues (with (34)), provided we accept to set accordingly the value of $d_{1}$, defined by (16), and hence the number $(\log_{2}d_{1}+1)$ of stages of the method and of eigendecompositions to be performed (and, again, provided this method remains numerically accurate enough, despite its higher computational complexity).
Moreover, also in this multi-stage extension, we aim at reducing the number $d_{2}$ of different eigenvalues, because we thus hope to better succeed in assigning each estimated eigenvalue to the correct value among the possible ones, despite estimation errors. A very attractive feature of the approach proposed here it that it allows one to decrease $d_{2}$ to 2, thus leading to Dichotomic multi-stage Eigenanalysis-based QPT, or DEQPT, methods (where “dichotomic” refers to the fact that the complete set of eigenvalues is split into two subsets, with identical values in each subset).
In Appendix K-A we prove that, for $d_{2}=2$, the above-defined subspace intersections are each indeed restricted to one dimension and yield all columns of $U$ (up to scale or phase factors). Moreover, we illustrate this property and thus the proposed complete multi-stage methods in Appendix K-B. Finally, Appendix K-C contains a pseudo-code for that multi-stage QPT method applicable to an arbitrary state space dimension and called EQPT5.
V Test results
To validate the EQPT1, EQPT2, EQPT3 and EQPT5 methods detailed in the previous sections, and to evaluate their accuracy, we performed numerical tests with data derived from a software simulation of the considered configuration. As mentioned in Section I, all the methods proposed in the present paper belong to the class of QPT methods that first perform a complete QST for each considered process output state and then exploit the results of these QST to achieve the core of QPT. The input of our QPT methods thus consists of estimates (more precisely, models of such estimates, as detailed below) of pure or mixed states provided by QST. Various QST methods were reported in the literature (see e.g. [1, 61, 62, 11, 9, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]) and one may hereafter e.g. use any of the available QST methods compatible with the types of quantum states encountered in each of the QPT methods proposed in the present paper. Therefore, to here provide numerical results that define the accuracy of our QPT methods, we do not restrict this investigation to the use of a single QST method, from which the reader could not easily extrapolate to his preferred QST approach, with his preferred parameter values (especially, the number of times he prepares a copy of a considered input state and he performs a measurement for the corresponding output state), whereas these parameter values may have a strong influence on QPT performance. Instead, we create a numerical model of quantum state estimates with a given accuracy and we focus on the resulting accuracy of our QPT methods, not only to estimate these absolute accuracies, but mainly to determine which of these methods yields the best performance when they use the same accuracy for QST. This moreover allows us to investigate the numerical performance of our QPT methods over a wide range of conditions, by varying the accuracy of the modeled quantum state estimates.
It should also be noted that some papers from the literature focus on reducing the complexity of QST algorithms (see e.g. [72, 73]). The overall complexity of a QST-based QPT algorithm therefore does not necessarily come from its QST part, as opposed to its subsequent part that consists of using QST results for the core of QPT. This also motivates us to perform tests as explained above, in order to uncouple the investigation of the complexity of the core of QPT (e.g. expressed in terms of required CPU time) from that of QST.
More precisely, to model a ket estimate (e.g. for the ket $\lvert\Psi_{2}\rangle$ of Section II-B), we compute its theoretical value in the considered conditions (e.g. using (6) with the input ket defined before (12) and a known process matrix $U$) and we add random, independent, complex-valued fluctuations to all components of this ket. Their real and imaginary parts are separately created with a random generator that has a probability density function that is uniform over the range $[-w/2,w/2]$, where the parameter $w$ defines the magnitude of the modeled estimation fluctuations (their variance is thus equal to $w^{2}/12$). This parameter is hereafter varied so as to investigate the sensitivity of the proposed QPT methods to the magnitude of QST estimation errors. This simple error model has the advantage of being easily interpretable. The error model for density matrix estimation is defined accordingly and is presented in Appendix L.
Each elementary test uses a randomly drawn $d\times d$ process matrix $U$, obtained as follows. We first create a $d\times d$ real-valued matrix with each element randomly, uniformly, drawn in the interval $[0,1]$. We then perform a QR decomposition of this matrix (e.g. using one of the algorithms defined in [48] p. 112, [74] p. 223, [75]). This especially yields the so-called “Q factor” of that QR decomposition (see e.g. the software package [75]). This factor is guaranteed to be a unitary matrix (see [48] p. 112, [74] p. 223, [75]) and we therefore use it as our matrix $U$.
Each above-defined matrix $U$ is used by feeding it with the process input states defined in the previous sections, computing the corresponding exact output states (e.g. using (1) and (6)), deriving their above-defined modeled estimates associated with QST and applying the considered QPT method to those estimates, so as to derive the process matrix estimate $\widehat{U}_{5}$.
The criterion used to evaluate the performance obtained in such a test aims at measuring how close we get to the case when $\widehat{U}_{5}$ is equal to $U$ up to a phase factor. The criterion used to this end is defined by first introducing
$NMSE(U,\widehat{U}_{5})=\min_{\theta}\left[\frac{1}{2d}||U-e^{i\theta}\widehat{U}_{5}||_{F}^{2}\right]$ (35)
where ||.$||.||_{F}$ stands for the Frobenius norm and where the phase $\theta$ is chosen so as to minimize this NMSE. The NMSE value resulting from this minimization may be shown to be equal to
$NMSE(U,\widehat{U}_{5})=\frac{1}{2d}\left[||U||_{F}^{2}+||\widehat{U}_{5}||_{F}^{2}-2\ |\mathrm{Tr}(U^{{\dagger}}\widehat{U}_{5})|\right].$ (36)
This criterion is a Normalized Mean squared Error (NMSE) since it is based on the sum of the squared moduli of the elements of the error matrix $(U-e^{i\theta}\widehat{U}_{5})$, then normalized by a scale factor chosen so that the maximum value of this parameter is equal to one when $U$ and $\widehat{U}_{5}$ are unitary (because their squared norms are then equal to $d$). The criterion that we eventually consider is the Normalized Root Mean Square Error (NRMSE) associated with the above NMSE:
$NRMSE(U,\widehat{U}_{5})=\sqrt{NMSE(U,\widehat{U}_{5})}.$ (37)
For each dimension $d$ of the state space, the results reported below were obtained by running 100 above-defined elementary tests, each with a different, randomly drawn, process matrix $U$ (except that we only performed 10 tests when applying the EQPT5 method to 11 qubits, as explained below). The performance figure then provided is the mean of $NRMSE(U,\widehat{U}_{5})$ over all these tests. This protocol is repeated for various values of $d$.
[…]
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=]FIG1.png Normalized Root Mean Square Error (NRMSE) of estimation of process matrix $U$ of proposed Quantum Process Tomography (QPT) methods EQPT1 (black solid lines), EQPT2 (blue dash doted lines) and EQPT3 (red dashed lines), vs. number $q$ of qubits, for several values of the parameter $w$ that defines the magnitude of the estimation errors of Quantum State Tomography (QST).
The reported tests were carried out with a standard PC (Intel(R) Core(TM) i5-7500 CPU running at 3.40 GHz with 16 GB of RAM), using the Matlab software environment. As a first step, Fig. V shows the results thus obtained with our single-stage and two-stage methods, namely EQPT1 (black solid lines), EQPT2 (blue dash doted lines) and EQPT3 (red dashed lines). The results for EQPT1 are provided for all values of the state space dimension $d$ that take the form $d=2^{q}$, where $q$ is the equivalent number of qubits, that is here varied from 2 to 13. $q=13$ is the maximum number that was accepted in the considered hardware and software environment without running out of memory (for non-optimized code). The same approach applies to EQPT2 and EQPT3 except that, due to (28), only even values of $q$ were considered. Moreover, for EQPT2 and EQPT3, performance turned out to often be somewhat degraded for the lowest possible state space dimension, namely $d=4$, that corresponds to $q=2$ qubits. This is not an issue, since EQPT2 and EQPT3 were developed to address higher values of $d$ (EQPT1 can be used instead for lower values of $d$). Therefore, for better readability, the results of EQPT2 and EQPT3 for $q=2$ are skipped in Fig. V. It should also be noted that the complexity of the configurations that could be tested was limited by memory requirements, as stated above, hence not by the computational load of the proposed algorithms, that remained reasonable up to $q=13$. For instance, for each of the 100 process matrix estimations, the average execution time used for a) creating data (thus only “modeling” QST instead of running a QST algorithm: see above), b) running the QPT algorithm and c) displaying results, was as follows, respectively for EQPT1, EQPT2 and EQPT3: 1) $\simeq 1$ ms, 2.6 ms and 2.8 ms for $q=4$ qubits, 2) 22 ms, 170 ms and 200 ms for $q=8$ qubits and 3) 23 s, 180 s and 230 s for $q=12$ qubits. The execution times for the other numbers of qubits are provided in Table I.
Method number of qubits
1 2 3 4 5 6 7 8 9 10 11 12 13
EQPT1 $\simeq 1$ ms $\simeq 1$ ms $\simeq 1$ ms $\simeq 1$ ms 2.2 ms 3.9 ms 7.2 ms 22 ms 68 ms 400 ms 3.1 s 23 s 170 s
EQPT2 2.6 ms 15 ms 170 ms 4.3 s 180 s
EQPT3 2.8 ms 18 ms 200 ms 5.0 s 230 s
EQPT5 2.3 ms 5.4 ms 27 ms 100 ms 580 ms 4.4 s 38 s 490 s 7600 s
$\simeq$ 8 mn $\simeq$ 2 h 07 mn
TABLE I: Execution times (averaged over tests with all values of $w$) of the EQPT1, EQPT2, EQPT3 and EQPT5 methods vs. number of qubits.
The following conlusions can be drawn from Fig. V, when comparing the proposed methods for any given QST quality, defined by the considered value of $w$. The two-stage methods EQPT2 and EQPT3 yield almost the same performance, with a slight advantage for EQPT3 (as expected from Appendix J), that is more significant for lower values of the space dimension $d$. Moreover, these two-stage methods always yield better and often much better performance than the single-stage method EQPT1, as expected from the beginning of Section III. Up to a certain space dimension $d$, the gap between these two types of methods increases with the space dimension and hence with the associated number $q$ of qubits: whereas all methods yield linear variations of the logarithm of NRMSE with respect to $q$ (i.e. logarithm of $d$), the slope of these variations is significantly lower for EQPT2 and EQPT3, and these methods already start with a lower NRMSE than EQPT1 for the lowest values of $d$. Depending on the considered conditions, the EQPT2 and EQPT3 methods decrease the NRMSE by a factor up to 6, and hence the NMSE by a factor up to 36, as compared with EQPT1. Moreover, for the lowest QST error magnitude considered in these tests, for all analyzed state space dimensions, EQPT2 and EQPT3 yield an NRMSE limited to about $10^{-1}$ and hence an NMSE limited to about $10^{-2}$.
[…]
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=]FIG2.png Normalized Root Mean Square Error (NRMSE) of estimation of process matrix $U$ of proposed Quantum Process Tomography (QPT) methods EQPT1 (black solid lines), EQPT2 (blue dash doted lines) and EQPT5 (red dashed lines), vs. (number $q$ of qubits) for several values of the parameter $w$ that defines the magnitude of the estimation errors of Quantum State Tomography (QST).
We now move to the multi-stage method EQPT5, whose performance is shown in Fig. I (red dashed lines), together with the results already provided above for EQPT1 (black solid lines) and EQPT2 (blue dash doted lines). The results for EQPT5 are reported for values of the state space dimension $d$ that take the form $d=2^{q}$, where $q$ is the equivalent number of qubits, varied as follows. For a large $q$, EQPT5 yields a significantly higher computational complexity than the single-stage and two-stage methods, due to the higher number of eigendecompositions it performs and to the associated program structure (recursive function with a binary tree structure, as explained in Appendix K). Unlike for the previous proposed methods, the applicability of EQPT5 to a large number of qubits was therefore more limited by its computational load than by its memory requirements, when using the considered basic hardware and software means. Therefore, to keep a reasonable simulation duration (for all estimations of the process matrix), EQPT5 was only tested up to $q=10$ qubits when performing 100 above-defined elementary tests, and also for $q=11$ qubits but then with only 10 elementary tests. Moreover, as with EQPT2, we hereafter skip the results for a low dimension, that is for $d\leq 4$, which is not the case of interest for a multi-stage method. In the considered range of space dimensions, the above-defined average execution time per process matrix estimation was e.g. as follows for EQPT5: 5.4 ms for $q=4$ qubits, 4.4 s for $q=8$ qubits, 490 s $\simeq$ 8 mn for $q=10$ qubits and 7600 s $\simeq$ 2 h 07 mn for $q=11$ qubits. The execution times for the other numbers of qubits are provided in Table I.
Fig. I yields the following conclusions for EQPT5. For the lowest considered values of the space dimension, EQPT5 starts with a somewhat higher NRMSE than EQPT1 and EQPT2. But, an attractive feature of EQPT5 is that its NRMSE then increases more slowly that those of the other methods, with respect to the space dimension. This again yields linear variations of the logarithm of NRMSE with respect to the number of qubits. EQPT5 thus outperforms the other two methods for higher space dimensions, and the gap between them increases with that dimension, up to a certain dimension. EQPT5 is thus attractive for high space dimensions, so that it actually reaches the goal for which it was designed. Depending on the considered conditions, EQPT5 decreases the NRMSE by a factor up to 3.2, and hence the NMSE by a factor up to more than 10, as compared with EQPT2. Moreover, for the lowest QST error magnitude considered in these tests, for all analyzed state space dimensions (which are slightly lower than those considered above for EQPT1 and EQPT2, however), EQPT5 yield an NRMSE limited to around $3.2\times 10^{-2}$ and hence an NMSE limited to around $10^{-3}$.
VI Conclusion and future works
Quantum Process Tomography (QPT) is a research topic of major interest, due to its importance for experimentally characterizing the actual behavior of quantum gates. In the present paper, we addressed QPT configurations that have two features: first, we consider unitary processes, and second, we aim at developing QPT methods that are applicable to a high state space dimension, corresponding to a significant number of qubits, as explained in Section I.
Since the considered process is unitary, its input-output relationship takes a specific form, that suggests that eigenanalysis may be used to estimate at least part of the parameters of the process matrix. We first built upon that idea, and we then extended it so as to also estimate the other process parameters that a plain eigenanalysis cannot estimate. This resulted in a first class of QPT methods (including EQPT1), that are stated to be single-stage since they only require one eigendecomposition.
The above approach may however be expected to yield limited performance when considering high-dimensional state spaces, due to the estimation errors of the Quantum State Tomography (QST) that it uses as a building block. We very significantly reduced this problem by introducing multi-stage eigenanalysis-based extensions of our approach, where the term “multi-stage” refers to the fact these methods successively perform several eigendecompositions. We explained in detail how two-stage methods (especially EQPT2 and EQPT3) operate, and then how to extend them to an arbitrary number of stages in order to minimize their sensitivity to QST errors (especially with the dichotomic EQPT5 method).
The relevance of this approach was validated by means of numerical tests performed with a simulated version of this QPT problem. This especially showed that, depending on the considered conditions, the tested two-stage methods (EQPT2 and EQPT3) decrease the NMSE by a factor up to 36, as compared with the tested single-stage method (EQPT1). Moreover, for the lowest QST error magnitude considered in these tests, for all analyzed state space dimensions, the two-stage methods yield an NMSE limited to about $10^{-2}$. Besides, the multi-stage dichotomic EQPT5 method may yield a higher accuracy (NMSE decreased by a factor up to 10 as compared with the two-stage EQPT2 method and NMSE limited to about $10^{-3}$ in almost the same conditions as above) but at the expense of a higher computational load.
We here had to restrict ourselves to tests performed on a standard PC. Our future works will therefore especially include additional tests, performed on a much more powerful platform, to derive the performance of our methods for a significantly higher number of qubits. Anyway, the configurations reported in the present paper already involve a quite significant number of qubits (up to 11 to 13, depending on the considered method) and hence a high state space dimension $d$ (up to a few thousands) as compared with various investigations from the literature, in particular with only $d=5$ in [3] that also deals with unitary processes. Our tests also showed that the applicability of our single-stage and two-stage methods was not limited by their complexity in terms of computational cost, unlike various alternative QPT methods, but only by the memory size available on the platform that was used for running them. Our dichotomic multi-stage methods have the opposite features.
Appendix A Connections between our QPT methods and classical signal processing
As stated in Section I, the type of QPT approaches proposed in this paper has connections with works that were previously reported in the classical domain. These connections may be defined as follows. QPT may be considered as the quantum version of so-called (classical) system identification [1]. The latter classical problem consists of estimating the parameter values of a transform from the input to the output of a system. That system may have multiple inputs, so that it then performs a combination of them. In the classical framework, such a combination is called a mixture. The classical problem then considered in the literature is therefore also referred to as mixture identification and it especially includes blind mixture identification, or BMI [59, 58, 57, 60]. This terminology should be used with care when also considering the quantum domain, because the above classical “mixtures” have nothing to do with the “statistical mixtures”, i.e. “mixed states”, faced in the quantum domain and used further in this paper. Only considering previous works in the classical domain at this stage of our discussion, the above-mentioned BMI has a close relationship with source separation, especially including blind source separation, or BSS [76, 77, 58, 57, 50, 78, 53]. BSS does not focus on identifying the transform performed by the above, i.e. direct, system itself, but mainly on deriving an estimate of the inverse of that transform, so as to apply it to the outputs of the above direct system and to thus restore the unknown source signals that form the input of the direct system. This relationship between BMI and BSS even appears in the name of the well-known SOBI method [79], which stands for “Second-Order Blind Identification”, whereas it is claimed to be a BSS method. The term “Second-Order” here refers to the fact that this method applies to random signals and uses the statistical properties of the measured signals, more precisely their usual second-order statistical parameters that consist of their covariance matrices involving different time delays between signals.
So far, we especially discussed how SOBI-like methods and the methods proposed in the present paper are connected in terms of the nature of data processing problems that they address. These two types of methods also have relationships in terms of the approaches that they use to solve the considered problems. More precisely, in classical BSS/BMI, when the transform performed by the direct system is linear (and memoryless) and its identification is performed by using the above-mentioned second-order statistics of the output signals of the direct system, an equation similar to (1) is obtained [80, 79, 60, 81, 82]. That classical counterpart of the structure of (1) was already exploited to perform classical BSS/BMI by means of several methods based on eigendecomposition. The real-valued version of the basic structure of (1) was first used in [80, 60] together with a required modified form of this equation (a related approach also appears in [82]). It was then extended in [79], by combining it with more than one modified form of this equation, to obtain a more robust method. Another version [81] aims at extending that type of approach from the standard case of stationary time-domain source signals to nonstationary time-domain sources. That version eventually uses complex-valued signals (whereas the mixing transform remains real-valued) because one moves to the Fourier domain.
Appendix B Creating a mixed state that has a diagonal density matrix
The algorithm proposed in Section II-A uses an input density matrix $\rho_{1}$ that is diagonal. It should be noted that, when one considers that one is able to create a pure state, then, creating the type of state considered here is not a problem, as will now be shown. We here consider the approach initially used by von Neumann in [83] to define a statistical mixture: such a state is expressed as being composed of a set of pure states, each of them being associated with a probability of occurrence. Then, to create experimental data corresponding to the state defined by the diagonal density matrix $\rho_{1}$, one just has to repeatedly, randomly, select one of the pure states that form the considered basis of the state space, moreover selecting the $k$th basis state with a probability that is equal to the $k$th value on the diagonal of $\rho_{1}$. This simplicity of implementation is worth noting because, in constrast, the authors of [44, 3] consider that using a mixed input state is too complex (maybe because they have non-diagonal density matrices in mind), so that they choose to restrict themselves to pure states, which has the drawback of requesting many such multiqubit input states, as opposed to only two multiqubit input states in the complete approach that we here start to describe (namely the mixed state with a density matrix $\rho_{1}$ considered here, and then the pure state $\lvert\Psi_{1}\rangle$ in Section II-B). Beyond their relationship based on considering (1), our approach and the one in [44] are therefore quite different.
Appendix C Preprocessing steps of the single-stage QPT algorithm
The first part of our single-stage algorithm, described in Section II-A, is based on an estimated density matrix $\widehat{\rho}_{2}$ provided by a QST algorithm. If this estimate $\widehat{\rho}_{2}$ is not guaranteed to be Hermitian due to the estimation errors of the considered “external” (i.e. external to our QPT) QST algorithm, a first recommended preprocessing step of our method consists of deriving a Hermitian estimate $\widehat{\rho}_{3}$ from $\widehat{\rho}_{2}$. This may be achieved by computing the Hermitian part of $\widehat{\rho}_{2}$, that is (see [48] p. 109 or [50] p. 343 for its real-valued counterpart):
$\widehat{\rho}_{3}=\frac{1}{2}\left(\widehat{\rho}_{2}+\widehat{\rho}_{2}^{{\dagger}}\right).$ (38)
Similarly, the actual matrix $\rho_{2}$ is known to have a trace equal to 1, since it is a density matrix (see e.g. [1] p. 101, [84] p. 73, [45] p. 283). Therefore, if its estimate ($\widehat{\rho}_{2}$ or) $\widehat{\rho}_{3}$ obtained so far is not guaranteed to have a unit trace due to QST estimation errors, a second recommended preprocessing step of our method consists of deriving a (still Hermitian and moreover) unit-trace estimate $\widehat{\rho}_{4}$ from ($\widehat{\rho}_{2}$ or) $\widehat{\rho}_{3}$, by dividing all its elements by its trace:
$\widehat{\rho}_{4}=\frac{\widehat{\rho}_{3}}{\mathrm{Trace}(\widehat{\rho}_{3})}.$ (39)
Similarly, the second part of our single-stage algorithm, described in Section II-B, uses an estimated ket $\lvert\widehat{\Psi}_{2}\rangle$, provided by a QST algorithm. If the considered QST algorithm does not guarantee that the norm of $\lvert\widehat{\Psi}_{2}\rangle$ is equal to 1, a recommended pre-processing step of our QPT method consists of dividing $\lvert\widehat{\Psi}_{2}\rangle$ by its norm.
Appendix D Preferred input density matrix for the first part of the single-stage QPT algorithm
In the first step of our first method defined in Section II-A, we preferably use the value of $\rho_{1}$ defined as follows. The values on its diagonal are the eigenvalues of the considered density operator $\rho_{2}$, they are real and non-negative and their sum is equal to one (see e.g. [1] p. 101). As shown in Section II-A, the proposed method assumes that these eigenvalues are different and ordered, and it is based on reordering the eigenvalues in the eigendecomposition of the estimated output density matrix and on permuting the eigenvectors accordingly. Therefore, if using quite close actual eigenvalues when choosing $\rho_{1}$, one may fear that the corresponding estimated eigenvalues obtained for the output density matrix be permuted due to estimation errors, so that the corresponding eigenvectors may then be permuted accordingly. This is a major problem, because the intermediate estimate $\widehat{U}_{2}$ related to the considered process, obtained at the output of this first step of our QPT method, is primarily based on these eigenvectors. In this paper, we handle this problem as follows: first, for the method imposed in Section II-A, we select the value of $\rho_{1}$ in order to limit that problem to the greatest extent that is possible within the frame of that method; then, in Sections III and IV, we propose more advanced methods that allow one to further reduce that problem. So, for the method considered here, we choose eigenvalues that are uniformly distributed (i.e. multiples of the same step, with zero excluded), so as to maximize the gap between any two adjacent values. Moreover constraining them to be positive, with a sum equal to one and in decreasing order as explained above, it is easily shown that one obtains the following values for the diagonal elements $\rho_{1,kk}$ of $\rho_{1}$:
$\rho_{1,kk}=\frac{2(d-k+1)}{d(d+1)}\hskip 14.22636ptk\in\{1,\dots,d\}.$ (40)
Appendix E Preferred input ket for the second part of the single-stage QPT algorithm
In the second step of first method defined in Section II-B, we select the input ket $\lvert\Psi_{1}\rangle$ as follows. Based on the condition on nonzero components provided in Section II-B for $\lvert\Psi_{1}\rangle$, we preferably select a state $\lvert\Psi_{1}\rangle$ with all component moduli as high as possible. Considering positive real-valued components and taking into account that $\lvert\Psi_{1}\rangle$ has unit norm, this yields a single solution: all components of $\lvert\Psi_{1}\rangle$ in the considered basis are equal to $1/\sqrt{d}$.
Appendix F Mathematical definitions and properties
Definition 1
[…]
Theorem 1
Every normal matrix $A\in M_{n}$ can be written as
$A=U\Lambda U^{{\dagger}},$ (41)
where $U\in M_{n}$ is unitary and $\Lambda\in M_{n}$ is a diagonal matrix whose main diagonal entries are the eigenvalues of $A$ ($M_{n}$ is the set of $n$-by-$n$ complex matrices).
Proof 1
See [48] p. 157 (e). See also [48] p. 101 for unitarily diagonalizable matrices (every normal matrix is unitarily diagonalizable) and the spectral theorem for normal matrices. In particular, this implies the spectral theorem for Hermitian matrices: see [48] p. 104.∎
Theorem 2
If $A\in M_{n}$ reads
$A=PDP^{-1}$ (42)
where $D\in M_{n}$ is a diagonal matrix and $P\in M_{n}$ is invertible then, for any index $k\in\{1,\dots,n\}$, the $k$th column of $P$ is an eigenvector of $A$ and the associated eigenvalue is the $k$th value on the diagonal of $D$.
Proof 2
Right-multiplying (42) by $P$ yields
$AP=PD.$ (43)
For any index $k\in\{1,\dots,n\}$, the $k$th column of the matrix equation (43) reads
$Av_{k}=\lambda_{k}v_{k}$ (44)
where $v_{k}$ is the $k$th column of $P$ and $\lambda_{k}$ is the $k$th element on the diagonal of $D$. Eq. (44) shows that $v_{k}$ is an eigenvector of $A$ and that the associated eigenvalue is $\lambda_{k}$.∎
Theorem 3
If $A\in M_{n}$ is normal, and if $x$ and $y$ are eigenvectors corresponding to distinct eigenvalues, $x$ and $y$ are orthogonal.
Proof 3
See [48] p. 103.∎
Theorem 4
A best least-squares approximation of a given matrix $A\in M_{n}$ by a unitary matrix $U_{A}\in M_{n}$ is given by $U_{A}=VW^{{\dagger}}$, where $A=V\Sigma W^{{\dagger}}$ is a singular value decomposition (SVD) of $A$.
Proof 4
See [48] p. 431.∎
Appendix G A variant for the first part of single-stage QPT methods
We here again consider the first part of single-stage QPT methods. As in the variant that we proposed in Section II-A for that first part, we here use a mixed input state, represented by the density matrix $\rho_{1}$. However, for the sake of generality, this matrix is here not any more constrained to be diagonal. It is still Hermitian and therefore unitarily diagonalizable. In the version considered here, as in Section II-A, we request all its eigenvalues to be different. That input density matrix $\rho_{1}$ is here requested to be known (i.e. fixed a priori or estimated from copies of that state, so that this variant is supervised or semi-supervised).
Using the same properties of eigendecomposition as in Section II-A, we consider all unitarily diagonalized forms of $\rho_{1}$ that have the following properties:
  1. 1.
    The eigenvalues are in a known order along the diagonal of $\rho_{1}$. Without loss of generality, we hereafter consider the case when they are in decreasing order.
  2. 2.
    All eigenvectors have unit norm.
All these eigendecompositions of $\rho_{1}$ may be expressed as
$\rho_{1}=U_{I1}\hskip 2.84526pt\rho_{I1D}\hskip 2.84526ptU_{I1}^{{\dagger}}$ (45)
where $U_{I1}$ and $\rho_{I1D}$ are respectively a unitary matrix and a diagonal matrix associated with that first input of the considered process.
Starting from a given $\rho_{1}$, the first step of the algorithm proposed in the present section consists of determining one (arbitrary) of these decompositions (45) meeting the above-defined conditions. This is performed by 1) running an eigendecomposition algorithm, then 2) reordering the values in the matrix obtained for $\rho_{I1D}$ if they are not directly provided in decreasing order by the considered eigendecomposition algorithm, 3) reordering the columns of the matrix obtained for $U_{I1}$ accordingly and 4) dividing these columns by their norms if the considered practical eigendecomposition algorithm is not guaranteed to yield unit-norm vectors. Thus, we especially get a known value $U_{I1}$ (which contains indeterminacies consisting of a phase factor separately for each of its columns). $U_{I1}$ is used in the second part of this version of our algorithm: see Appendix I.
Besides, using (45), Equation (1) may be expressed as
$\rho_{2}=U_{g}\hskip 2.84526pt\rho_{I1D}\hskip 2.84526ptU_{g}^{{\dagger}}$ (46)
where the global unitary matrix
$U_{g}=UU_{I1}$ (47)
includes the effect of the eigendecomposition of $\rho_{1}$, in addition to the considered process $U$.
Formally, (46) adresses exactly the same problem as in Section II-A (see (1) with a diagonal $\rho_{1}$), but now with a diagonal input density matrix $\rho_{I1D}$ and a process $U_{g}$ (that is also unitary). We can therefore apply the algorithm and results of Section II-A to this reformulated problem, i.e., briefly, we diagonalize $\widehat{\rho}_{4}$ and we then reorder the eigenvalues (in the same order as in $\rho_{I1D}$) and associated normalized eigenvectors. This guarantees that the algorithm of Section II-A here provides an estimate $\widehat{U}_{2}$ that restores $U_{g}$ up to a phase factor for each of its columns, with (5) here replaced by
$\widehat{U}_{2}$ $=$ $U_{g}D_{\phi}$ (48)
$=$ $U\hskip 2.84526ptU_{I1}\hskip 2.84526ptD_{\phi}$ (49)
where $D_{\phi}$ is again a diagonal matrix that contains unknown phase factors $e^{\phi_{1}},\dots,e^{\phi_{d}}$ on its main diagonal.
As opposed to the basic version of the first step defined in Section II-A, the estimate $\widehat{U}_{2}$ here contains the effect of $U_{I1}$ in addition. The second part of the method must then be modified accordingly, as compared with its basic version defined in Section II-B. A solution to this problem is provided in Appendix I.
Appendix H A variant for the second part of single-stage QPT methods
In the framework of single-stage QPT methods that include a first part as defined in Section II-A, we here again consider the associated second part, that aims at estimating the phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$, by using a second input state of the considered process. The method that we proposed in Section II-B for that second part uses a pure input state defined by a ket $\lvert\Psi_{1}\rangle$. However, a mixed input state having a density matrix $\rho_{5}$ may be used instead, as will now be shown. As $\lvert\Psi_{1}\rangle$ in Section II-B, the input density matrix $\rho_{5}$ is here supposed to be initially known (supervised case) or estimated by using part of the available copies of that state (semi-supervised case).
The input state $\rho_{5}$ yields a state $\rho_{6}$ at the output of the considered process, and these two states have the same relationship as in (1). In practice, what is available, thanks to QST, is an estimate $\widehat{\rho}_{8}$ of $\rho_{6}$, here again after it has been pre-processed, if required, so as to be Hermitian and to have unit trace. Thus, up to estimation errors, $\widehat{\rho}_{8}$ meets the condition
$\widehat{\rho}_{8}=U\hskip 2.84526pt\rho_{5}\hskip 2.84526ptU^{{\dagger}}.$ (50)
Then, knowing $\widehat{\rho}_{8}$ and $\widehat{U}_{2}$ obtained in the first part of our algorithm (see Section II-A), the next step of our algorithm here consists of combining them to form
[…]
This may be seen as the extension, to a mixed state, of the quantity (9) in the variant of the present method that uses a pure input state instead: for the latter type of state, $\widehat{\rho}_{8}$ reduces to
$\widehat{\rho}_{8}\ =\lvert\widehat{\Psi}_{2}\rangle\langle\widehat{\Psi}_{2}\lvert$ (52)
and the state $\lvert\Psi_{3}\rangle$ considered in Section II-A corresponds to $\rho_{9}=\lvert\Psi_{3}\rangle\langle\Psi_{3}\lvert$. Combining the latter expression with (9) and using (52), the resulting expression of $\rho_{9}$ is fully compatible with its extension (51) to an arbitrary mixed input state considered here.
Combining (50), (51) and (5), with $U$ unitary, yields
$\rho_{9}=D_{\phi}^{*}\hskip 2.84526pt\rho_{5}\hskip 2.84526ptD_{\phi}.$ (53)
This shows that a diagonal matrix $\rho_{5}$ cannot be used in this approach, because all three matrices in the right-hand term of (53) are then diagonal, so that $D_{\phi}^{*}$ and $D_{\phi}$ cancel out and (53) reduces to
$\rho_{9}=\rho_{5}$ (54)
and $\rho_{9}$ does not yield any information about the phase factors in $D_{\phi}$. This is not surprising because all the information that can be extracted with a diagonal input density matrix was already extracted by using such a matrix in the first part of our algorithm (see Section II-A). We therefore consider a non-diagonal matrix $\rho_{5}$ hereafter.
Whatever the input density matrix $\rho_{5}$, denoting $\rho_{kl}$ its elements, (53) yields (again up to estimation errors)
$\rho_{9}=\left[\begin{tabular}[]{llll}$\rho_{11}$&$\rho_{12}e^{i(\phi_{2}-\phi_{1})}$&$\dots$&$\rho_{1d}e^{i(\phi_{d}-\phi_{1})}$\\ $\dots$&&&$\dots$\end{tabular}\right].$ (55)
Therefore, we consider a value of $\rho_{5}$ that is arbitrary except that all the elements of its first row are nonzero (this method may be adapted to the case when another row only has nonzero elements). The next step of our algorithm consists of computing the quantity
$\left[\rho_{9}\right]_{1\bullet}\oslash\left[\rho_{5}\right]_{1\bullet}$ (56)
where [.$\left[.\right]_{1\bullet}$ represents the row vector equal to the first row of the considered matrix (i.e. row 1 and all columns, or “columns $\bullet$”). Thanks to (55), we have
$\left[\rho_{9}\right]_{1\bullet}\oslash\left[\rho_{5}\right]_{1\bullet}=\left[1,e^{i(\phi_{2}-\phi_{1})},\dots,e^{i(\phi_{d}-\phi_{1})}\right]$ (57)
and hence
$\mathrm{diag}((\left[\rho_{9}\right]_{1\bullet}\oslash\left[\rho_{5}\right]_{1\bullet})^{*})=e^{i\phi_{1}}D_{\phi}^{*}.$ (58)
The next step of our algorithm therefore consists of computing
$\widehat{U}_{5}$ $=$ $\widehat{U}_{2}\hskip 2.84526pt\mathrm{diag}((\left[\rho_{9}\right]_{1\bullet}\oslash\left[\rho_{5}\right]_{1\bullet})^{*})$ (59)
$=$ $\widehat{U}_{2}\hskip 2.84526pt\mathrm{diag}((\left[\widehat{U}_{2}^{{\dagger}}\hskip 2.84526pt\widehat{\rho}_{8}\hskip 2.84526pt\widehat{U}_{2}\right]_{1\bullet}\oslash\left[\rho_{5}\right]_{1\bullet})^{*})$ (60)
where (60) results from (51) and defines the complete processing performed in this second part of our algorithm, starting from the estimate $\widehat{U}_{2}$ of $U$ obtained in the first part of the algorithm and combining it with the density matrices $\rho_{5}$ and $\widehat{\rho}_{8}$ used in that second part.
Combining (59), (58) and (5) yields
$\widehat{U}_{5}=e^{i\phi_{1}}U$ (61)
i.e. $\widehat{U}_{5}$ succeeds in restoring $U$ (once again, up to a global phase factor that has no physical consequences and up to the estimation errors that were mentioned above but implicit in the above equations).
Appendix I Another variant for the second part of single-stage QPT methods
In the framework of single-stage QPT methods, we here again consider the case when the first part is defined as in Appendix G. We present an associated solution for the second part of the method, using a pure input state defined by a ket $\lvert\Psi_{1}\rangle$. $\lvert\Psi_{1}\rangle$ is requested to be known, i.e. fixed a priori (supervised case) or estimated from a subset of its copies (semi-supervised case).
Starting with the same approach as in Section II-B, this here again yields the output ket estimate (8). Our algorithm then computes $\lvert\Psi_{3}\rangle$ defined by (9), as in Section II-B, but now with $\widehat{U}_{2}$ that meets (49). Combining (49) with (9) and (8) and considering that $U$ is unitary (hence $U^{{\dagger}}U=I$) , we get
$\lvert\Psi_{3}\rangle=e^{i\theta}D_{\phi}^{*}\lvert\Psi_{4}\rangle$ (62)
where we introduce
$\lvert\Psi_{4}\rangle=U_{I1}^{{\dagger}}\lvert\Psi_{1}\rangle$ (63)
which is a known quantity. Comparing (62) with (10) shows that $\lvert\Psi_{4}\rangle$ here replaces the ket $\lvert\Psi_{1}\rangle$ used in the method of Section II-B (which is coherent with the fact that the latter method is a reduced form of the present one, for the case when $\rho_{1}$ is diagonal and hence $U_{I1}$ is the identity matrix). The remainder of the present method is then derived by modifying the end of the method of Section II-B accordingly: here using
$\widehat{U}_{5}=\widehat{U}_{2}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{4}\rangle)U_{I1}^{{\dagger}}$ (64)
it may be shown in the same way as in Section II-B that we here again get (14), i.e. $\widehat{U}_{5}$ here again succeeds in restoring $U$, up to a global phase factor (and again up to estimation errors).
Appendix J Variants of two-stage QPT methods
Since the columns of $\widehat{U}_{4}$ are estimated independently, one may fear that they are not exactly orthogonal, due to estimation errors. Then, $\widehat{U}_{4}$ and hence $\widehat{U}_{5}$ are not unitary, although they should be because they aim at estimating $U$, that is constrained to be unitary. A first solution to this problem consists of postprocessing $\widehat{U}_{4}$, so as to determine a best least-squares unitary approximation of $\widehat{U}_{4}$. An algorithm that may be used to this end is defined by Theorem 4. The resulting modified QPT method is called EQPT3 hereafter. Its pseudo-code is provided in Algorithm 3. A slightly different version consists of finding a unitary approximation of $\widehat{U}_{5}$ instead of $\widehat{U}_{4}$. The resulting QPT method is called EQPT4. Its pseudo-code is provided in Algorithm 4.
Input : a) Estimate $\widehat{\rho}_{4}$ of output density matrix $\rho_{2}$ (provided by QST and obtained for input density matrix (17) with the diagonal values defined in (19)). b) Estimate $\widehat{\rho}_{8}$ of output density matrix $\rho_{6}$ (provided by QST and obtained for input density matrix (21) with the diagonal values defined in (19)). c) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
Output : Estimate $\widehat{U}_{5}$ of quantum process matrix $U$.
begin
       /* Exploit $\widehat{\rho}_{4}$: */
       Eigendecomposition: derive 1) a diagonal matrix $D_{4}$ that contains the eigenvalues of $\widehat{\rho}_{4}$ in an arbitrary order and 2) a matrix $V_{4}$ whose columns are eigenvectors of $\widehat{\rho}_{4}$ in the same order as eigenvalues;
       Reorder the eigenvalues in $D_{4}$ in nonincreasing order and apply the same permutation to the columns of $V_{4}$ to create the matrix $\widehat{U}_{2}$;
       /* Exploit $\widehat{\rho}_{8}$: */
       Eigendecomposition: derive 1) a diagonal matrix $D_{8}$ that contains the eigenvalues of $\widehat{\rho}_{8}$ in an arbitrary order and 2) a matrix $V_{8}$ whose columns are eigenvectors of $\widehat{\rho}_{8}$ in the same order as eigenvalues;
       Reorder the eigenvalues in $D_{8}$ in nonincreasing order and apply the same permutation to the columns of $V_{8}$ to create the matrix $\widehat{U}_{3}$;
       /* Combine $\widehat{U}_{2}$ and $\widehat{U}_{3}$: */
       for $m_{1}=1$ to $d_{2}$ do
             for $m_{2}=1$ to $d_{2}$ do
                   Set column of $\widehat{U}_{4}$ with index equal to $((m_{1}-1)d_{1}+m_{2})$: Set it to first vector provided by CCA of 1) the set of columns of $\widehat{U}_{2}$ with indices $((m_{1}-1)d_{1}+n_{1})$ with $n_{1}\in\{1,\dots,d_{1}\}$ and 2) the set of columns of $\widehat{U}_{3}$ with indices $((m_{2}-1)d_{1}+n_{2})$ with $n_{2}\in\{1,\dots,d_{1}\}$;
                   /* If the CCA tool does not provide unit-norm vectors, then do it: */
                   Divide above column of $\widehat{U}_{4}$ by its norm;
             end for
       end for
      Find decomposition of $\widehat{U}_{4}$ as $\widehat{U}_{4}=V\Sigma W^{{\dagger}}$ with SVD;
       $\widehat{U}_{4}{}_{u}=VW^{{\dagger}}$;
       /* Exploit $\lvert\widehat{\Psi}_{2}\rangle$: */
       $\lvert\Psi_{3}\rangle=\widehat{U}_{4}{}_{u}\hskip 0.28453pt^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}=\widehat{U}_{4}{}_{u}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
end
Algorithm 3 EQPT3: third Eigenanalysis-based QPT algorithm (composed of two stages). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see text.
Input : a) Estimate $\widehat{\rho}_{4}$ of output density matrix $\rho_{2}$ (provided by QST and obtained for input density matrix (17) with the diagonal values defined in (19)). b) Estimate $\widehat{\rho}_{8}$ of output density matrix $\rho_{6}$ (provided by QST and obtained for input density matrix (21) with the diagonal values defined in (19)). c) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
Output : Estimate $\widehat{U}_{5}$ of quantum process matrix $U$.
begin
       /* Exploit $\widehat{\rho}_{4}$: */
       Eigendecomposition: derive 1) a diagonal matrix $D_{4}$ that contains the eigenvalues of $\widehat{\rho}_{4}$ in an arbitrary order and 2) a matrix $V_{4}$ whose columns are eigenvectors of $\widehat{\rho}_{4}$ in the same order as eigenvalues;
       Reorder the eigenvalues in $D_{4}$ in nonincreasing order and apply the same permutation to the columns of $V_{4}$ to create the matrix $\widehat{U}_{2}$;
       /* Exploit $\widehat{\rho}_{8}$: */
       Eigendecomposition: derive 1) a diagonal matrix $D_{8}$ that contains the eigenvalues of $\widehat{\rho}_{8}$ in an arbitrary order and 2) a matrix $V_{8}$ whose columns are eigenvectors of $\widehat{\rho}_{8}$ in the same order as eigenvalues;
       Reorder the eigenvalues in $D_{8}$ in nonincreasing order and apply the same permutation to the columns of $V_{8}$ to create the matrix $\widehat{U}_{3}$;
       /* Combine $\widehat{U}_{2}$ and $\widehat{U}_{3}$: */
       for $m_{1}=1$ to $d_{2}$ do
             for $m_{2}=1$ to $d_{2}$ do
                   Set column of $\widehat{U}_{4}$ with index equal to $((m_{1}-1)d_{1}+m_{2})$: Set it to first vector provided by CCA of 1) the set of columns of $\widehat{U}_{2}$ with indices $((m_{1}-1)d_{1}+n_{1})$ with $n_{1}\in\{1,\dots,d_{1}\}$ and 2) the set of columns of $\widehat{U}_{3}$ with indices $((m_{2}-1)d_{1}+n_{2})$ with $n_{2}\in\{1,\dots,d_{1}\}$;
                   /* If the CCA tool does not provide unit-norm vectors, then do it: */
                   Divide above column of $\widehat{U}_{4}$ by its norm;
             end for
       end for
      /* Exploit $\lvert\widehat{\Psi}_{2}\rangle$: */
       $\lvert\Psi_{3}\rangle=\widehat{U}_{4}^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}{}_{p}\hskip 0.28453pt=\widehat{U}_{4}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
       Find decomposition of $\widehat{U}_{5}{}_{p}$ as $\widehat{U}_{5}{}_{p}=V\Sigma W^{{\dagger}}$ with SVD;
       $\widehat{U}_{5}=VW^{{\dagger}}$;
end
Algorithm 4 EQPT4: fourth Eigenanalysis-based QPT algorithm (composed of two stages). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see text.
Appendix K Additional information for the proposed multi-stage methods
K-A Relevance of the methods
In this appendix, we first prove that the approach with $d_{2}=2$ that we proposed in Section IV, based on subspace intersections, is relevant.
Each subspace ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$ defined in Section IV is associated with a set of columns of the process matrix $U$ to be identified, in the following sense. We consider a given arbitrary stage of our algorithm, defined by the index ${b_{\ell}}$, and the subset of columns of $U$ that correspond to (i.e. that have the same column indices as) a given arbitrary diagonal value $r_{m_{{b_{\ell}}}}$ of the input density matrix $\rho_{1}$ when forming (1). This subset of columns of $U$ and $r_{m_{{b_{\ell}}}}$ are defined by the index $m_{{b_{\ell}}}\in\{1,2\}$. We then have the following property: the eigendecomposition plus reordering algorithm proposed in Section IV yields a set of column vectors associated with ${b_{\ell}}$ and $m_{{b_{\ell}}}$, that form a basis of a subspace denoted as ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$, and that subspace is also the subspace spanned by the columns of $U$ associated with the index $m_{{b_{\ell}}}$ in the above-defined sense.
The analysis of the indices of these columns presented in the present appendix is made easier by introducing the following modified notations, related to the expression of these indices in base 2, which is motivated by the structure of the matrix $\rho_{1}$ based on powers of 2 that was introduced in Section IV:
  • The index of the first row and first column of $\rho_{1}$ and $U$ is here equal to 0 (instead of 1 in the other parts of this paper).
  • Each variable $m_{{b_{\ell}}}$, used to split in two subsets the columns of $\rho_{1}$, takes the values 1 and 2 (because here $d_{2}=2$). With this, we here associate another binary variable, that takes the values 0 and 1, namely $\mu_{{b_{\ell}}}=m_{{b_{\ell}}}-1$.
  • When using the variable $\mu_{{b_{\ell}}}$, the subspaces are denoted with the letter ${\cal T}$, to avoid any misunderstanding with the notation ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$ used above, especially when considering explicit numerical values of their indices: for instance, ${\cal T}_{0,0}$ is nothing but ${\cal S}_{0,1}$. More generally speaking, we have ${\cal T}_{{b_{\ell}},\mu_{{b_{\ell}}}}={\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$, whatever the values of the indices.
With these notations, using the description provided in Section IV, the indices of the columns of $U$ corresponding to ${\cal T}_{{b_{\ell}},\mu_{{b_{\ell}}}}$ may be shown to be
$2b_{s}b_{i}+\mu_{{b_{\ell}}}b_{s}+\{0,\dots,(b_{s}-1)\}$
$\hskip 56.9055pt\mathrm{with}\hskip 14.22636ptb_{i}\in\{0,\dots,(b_{n}-1)\}.$ (65)
This is more easily interpreted by indexing the stages of the algorithm in the reverse order, thus introducing
${b_{r}}={b_{\ell m}}-{b_{\ell}}$ (66)
where ${b_{\ell m}}=\log_{2}d_{1}$ is the maximum value of ${b_{\ell}}$ and therefore
$d_{1}=2^{b_{\ell m}}.$ (67)
Using (30), (31), (67) and (66), we obtain
$b_{s}=2^{b_{r}}.$ (68)
Also using (31), the column indices in (65) may therefore be rewritten as
$b_{i}2^{({b_{r}}+1)}+\mu_{{b_{\ell}}}2^{b_{r}}+\{0,\dots,(2^{b_{r}}-1)\}$
$\hskip 56.9055pt\mathrm{with}\hskip 14.22636ptb_{i}\in\{0,\dots,(2^{b_{\ell}}-1)\}.$ (69)
This should be compared to the representation in base 2 of the index of any column of $U$, that may be expressed as
$\sum_{j=0}^{{b_{\ell m}}}c_{j}2^{j}$ (70)
with each bit $c_{j}$ equal to 0 or 1. Eq. (70) allows one to interpret the set of indices in (69), i.e. the set of indices of the columns of $U$ that correspond to ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}={\cal T}_{{b_{\ell}},\mu_{{b_{\ell}}}}$, for given values of $m_{{b_{\ell}}}$, ${b_{\ell}}$ and hence ${b_{r}}$, due to (66): these column indices are all the indices whose representation in base 2 is such that their bit $c_{j}$ with index $j={b_{r}}$ is constrained to be equal to $\mu_{{b_{\ell}}}$, whereas all possible combinations are gathered in this set of index values, for all the bits $c_{j}$ with indices $j\neq{b_{r}}$. To summarize, in ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$, only the bit with index $j={b_{r}}$ of the column index is imposed.
Then, we eventually aim at analyzing the intersections of $({b_{\ell m}}+1)$ subspaces introduced in Section IV. This may be rephrased as follows, using the notations introduced in the present appendix. Each considered intersection of $({b_{\ell m}}+1)$ subspaces ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}={\cal T}_{{b_{\ell}},\mu_{{b_{\ell}}}}$ corresponds to one value of the set of indices $\{m_{0},\dots,m_{{b_{\ell m}}}\}$ and therefore $\{\mu_{0},\dots,\mu_{{b_{\ell m}}}\}$. As explained above, separately for each value of ${b_{\ell}}$ and hence of ${b_{r}}$, the columns of $U$ that are kept in this intersection are those whose index represented in base 2 has a bit with index $j={b_{r}}$ that is fixed and equal to $\mu_{{b_{\ell}}}$. The intersection of the complete set of $({b_{\ell m}}+1)$ subspaces obtained when varying ${b_{\ell}}$ thus yields one and only one column of $U$, namely the one whose bits are all defined by the selected complete set of values $\{\mu_{0},\dots,\mu_{{b_{\ell m}}}\}$. This also entails that, when successively considering the subspace intersections corresponding to all values of $\{\mu_{0},\dots,\mu_{{b_{\ell m}}}\}$, all columns of $U$ are successively obtained. This proves that the proposed method based on subspace intersections is guaranteed to completely identify $U$ (here again up to one phase factor per column, these phase factors being then handled as in the other proposed methods).
The successive subspace intersections may be considered in such an order that the sets of indices $\{\mu_{0},\dots,\mu_{{b_{\ell m}}}\}$ correspond to the representation in base 2 of the integers in increasing order. Then, the columns of $U$ are obtained with indices in increasing order. This is exactly what is obtained in the successive rows of Table II in the example provided below.
K-B Example
We here illustrate the above property and thus the proposed complete multi-stage methods by considering the case $d=8$, $d_{2}=2$ and hence $d_{1}=4$ and $(\log_{2}d_{1}+1)=3$ stages. During the first stage, that corresponds to ${b_{\ell}}=0$, the eigendecomposition allows one to identify two subspaces ${\cal S}_{0,m_{0}}$, respectively with $m_{0}=1$ and $m_{0}=2$, and that respectively correspond to the following sets of 4 indices of columns of $U$: $\{1,2,3,4\}$ and $\{5,6,7,8\}$. Then, during the second stage, that corresponds to ${b_{\ell}}=1$, the eigendecomposition allows one to identify two subspaces ${\cal S}_{1,m_{1}}$, respectively with $m_{1}=1$ and $m_{1}=2$, and that respectively correspond to the following sets of 4 indices of columns of $U$: $\{1,2,5,6\}$ and $\{3,4,7,8\}$. Finally, the third stage, that corresponds to ${b_{\ell}}=2$, yields the following sets of 4 indices of columns of $U$: $\{1,3,5,7\}$ and $\{2,4,6,8\}$. We then derive the corresponding subspace intersections, first considering ${\cal S}_{0,m_{0}}\cap{\cal S}_{1,m_{1}}$ as intermediate quantities, and finally deriving ${\cal S}_{0,m_{0}}\cap{\cal S}_{1,m_{1}}\cap{\cal S}_{2,m_{2}}$. All these intersections are listed in Table II. This proves that the proposed approach yields 8 intersections that are each restricted to one column of $U$ and that this set of intersections yields all the individual columns of $U$, as required.
${\cal S}_{0,m_{0}}$ ${\cal S}_{1,m_{1}}$ ${\cal S}_{2,m_{2}}$ ${\cal S}_{0,m_{0}}\cap{\cal S}_{1,m_{1}}$ ${\cal S}_{0,m_{0}}\cap{\cal S}_{1,m_{1}}\cap{\cal S}_{2,m_{2}}$
$\{1,2,3,4\}$$\{1,2,5,6\}$$\{1,3,5,7\}$$\{1,2\}$$\{1\}$
$\{1,2,3,4\}$$\{1,2,5,6\}$$\{2,4,6,8\}$$\{1,2\}$$\{2\}$
$\{1,2,3,4\}$$\{3,4,7,8\}$$\{1,3,5,7\}$$\{3,4\}$$\{3\}$
$\{1,2,3,4\}$$\{3,4,7,8\}$$\{2,4,6,8\}$$\{3,4\}$$\{4\}$
$\{5,6,7,8\}$$\{1,2,5,6\}$$\{1,3,5,7\}$$\{5,6\}$$\{5\}$
$\{5,6,7,8\}$$\{1,2,5,6\}$$\{2,4,6,8\}$$\{5,6\}$$\{6\}$
$\{5,6,7,8\}$$\{3,4,7,8\}$$\{1,3,5,7\}$$\{7,8\}$$\{7\}$
$\{5,6,7,8\}$$\{3,4,7,8\}$$\{2,4,6,8\}$$\{7,8\}$$\{8\}$
TABLE II: Multi-stage QPT method. Each column of the table contains the indices of the column vectors that correspond to the considered subspace or subspace intersection defined in the first row of the table.
K-C Pseudo-code
The above analysis also has the advantage of suggesting a possible structure for the software program used to implement this multi-stage QPT method. This program has a recursive structure, combined with a binary tree as follows. It is based on a recursive function, with each call of this function corresponding to one value of the stage index ${b_{\ell}}$ (and hence of ${b_{r}}$) and to one bit of the binary representation (70) of the index of a column of $U$. In each such call, this function successively considers the two possible values of that bit, that is 0 and 1. For each of them, it determines the intersection of the subspaces with indices 0 to ${b_{\ell}}$ (by determining the subspace with index ${b_{\ell}}$ and combining it with the previously determined intersection of the subspaces with indices 0 to $({b_{\ell}}-1)$) and it recursively calls itself with this bit value. Successively performing this for each of the two values of the considered bit creates two branches in the tree structure. In each final call of this recursive function, i.e. each leave of the tree, all bits of the column index are assigned, the function determines the corresponding complete subspace intersection and sends it back as one column of the estimate of $U$. By gathering these columns at each level back from the recursion, the complete matrix that estimates $U$ (up to phase factors at this stage) is obtained.
The corresponding pseudo-code is provided in Algorithm 5 for the main part of the program and Algorithm 6 for the recursive function.
Input : a) Set $\{\widehat{\rho}_{out,{b_{\ell}}}\}$ of estimates of output density matrices of process (provided by QST and obtained for input density matrices defined in Section IV): one estimate for each stage ${b_{\ell}}$ of recursion in Algorithm 6, b) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
Output : Estimate $\widehat{U}_{5}$ of quantum process matrix $U$.
begin
       /* Exploit $\{\widehat{\rho}_{out,{b_{\ell}}}\}$: */
       $\widehat{U}_{2}$= EQPT5-sub(0, $\{\widehat{\rho}_{out,{b_{\ell}}}\}$,0);
       /* Exploit $\lvert\widehat{\Psi}_{2}\rangle$: */
       $\lvert\Psi_{3}\rangle=\widehat{U}_{2}^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}=\widehat{U}_{2}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
end
Algorithm 5 EQPT5-main: main part of fifth Eigenanalysis-based QPT algorithm (composed of multiple stages). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see text.
Input : a) Index ${b_{\ell}}$ of recursion stage, b) Set $\{\widehat{\rho}_{out,{b_{\ell}}}\}$ of estimates of output density matrices of process (see Algorithm 5), c) Matrix ${\cal I}$ whose column vectors form a basis of subspace intersection determined in previous stages of recursion (none in first stage).
Output : matrix ${\widehat{U}_{part}}$ related to estimation of part of quantum process matrix $U$.
begin
      /* Set auxiliary variables (using $d$: state space dimension; $d_{1}=d/2$): */
       ${b_{\ell m}}=\log_{2}d_{1}$;
       $b_{n}=2^{b_{\ell}}$;
       $b_{s}=d_{1}/b_{n}$;
       /* Exploit single estimated output density matrix $\widehat{\rho}_{out,{b_{\ell}}}$ corresponding to current stage ${b_{\ell}}$ of recursion: */
       Eigendecomposition: derive 1) a diagonal matrix $D_{4}$ that contains the eigenvalues of $\widehat{\rho}_{out,{b_{\ell}}}$ in an arbitrary order and 2) a matrix $V_{4}$ whose columns are eigenvectors of $\widehat{\rho}_{out,{b_{\ell}}}$ in the same order as eigenvalues;
       Reorder the eigenvalues in $D_{4}$ in nonincreasing order and apply the same permutation to the columns of $V_{4}$ to create the matrix $\widehat{U}_{{b_{\ell}}}$;
       Store the first $d_{1}$ columns and last $d_{1}$ columns of $\widehat{U}_{{b_{\ell}}}$ respectively in matrices $\widehat{U}_{{b_{\ell}}}(1)$ and $\widehat{U}_{{b_{\ell}}}(2)$;
       /* Derive intersections ${\cal I}(1)$ and ${\cal I}(2)$ of subspaces defined by $\widehat{U}_{{b_{\ell}}}(1)$ and $\widehat{U}_{{b_{\ell}}}(2)$ with subspace defined by ${\cal I}$ (if any): */
       if ${b_{\ell}}=0$ then
             ${\cal I}(1)=\widehat{U}_{{b_{\ell}}}(1)$;
             ${\cal I}(2)=\widehat{U}_{{b_{\ell}}}(2)$;
       end if
      else
            ${\cal I}(1)=$ first $b_{s}$ column vectors provided by CCA of ${\cal I}$ and $\widehat{U}_{{b_{\ell}}}(1)$;
             ${\cal I}(2)=$ first $b_{s}$ column vectors provided by CCA of ${\cal I}$ and $\widehat{U}_{{b_{\ell}}}(2)$;
       end if
      if ${b_{\ell}}<{b_{\ell m}}$ then
             $\widehat{U}_{part}(1)$ = EQPT5-sub $({b_{\ell}}+1,\{\widehat{\rho}_{out,{b_{\ell}}}\},{\cal I}(1))$ ;
             $\widehat{U}_{part}(2)$ = EQPT5-sub $({b_{\ell}}+1,\{\widehat{\rho}_{out,{b_{\ell}}}\},{\cal I}(2))$ ;
       end if
      else
             /* If the CCA tool does not provide unit-norm vectors, then do it: */
             $\widehat{U}_{part}(1)={\cal I}(1)/\mathrm{norm}({\cal I}(1))$;
             $\widehat{U}_{part}(2)={\cal I}(2)/\mathrm{norm}({\cal I}(2))$;
       end if
      $\widehat{U}_{part}$ = concatenation of columns of $\widehat{U}_{part}(1)$ and $\widehat{U}_{part}(2)$;
end
Algorithm 6 EQPT5-sub: recursively called sub-program of fifth Eigenanalysis-based QPT algorithm (composed of multiple stages). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see text.
Appendix L Fluctuation model for density matrix estimation
In Section V, we presented the statistical model that we used in our tests to represent the fluctuations that may be faced when estimating a ket with a QST algorithm. We hereafter describe a corresponding model for the estimation of density matrices. Whereas we eventually apply it to mixed states, we first build it by considering a pure state, to make it compatible with the model used for ket estimation in Section V. We therefore consider a pure state represented by the density matrix
$\rho=\lvert\psi\rangle\langle\psi\rvert.$ (71)
Denoting $c_{k}$ each actual component of $\lvert\psi\rangle$ and $\epsilon_{k}$ the additive random complex-valued error made when estimating it, each actual element of the density matrix $\rho$ reads
$\rho_{k\ell}=c_{k}c_{\ell}^{*}$ (72)
and its estimated version may be expressed as
$\widehat{\rho}_{k\ell}$ $=$ $(c_{k}+\epsilon_{k})(c_{\ell}+\epsilon_{\ell})^{*}$ (73)
$=$ $c_{k}c_{\ell}^{*}+c_{k}\epsilon_{\ell}^{*}+\epsilon_{k}c_{\ell}^{*}+\epsilon_{k}\epsilon_{\ell}^{*}.$ (74)
We aim at building an approximation of that quantity that, as with the “noisy ket model” of Section V, yields a noisy model of the density matrix that is expressed as the sum of the actual density matrix and of a term that represents fluctuations that are consistent with those of the random model used in Section V for kets. To this end, we first note that, in (74), the first term, namely $c_{k}c_{\ell}^{*}$, is nothing but the actual value (72), whereas all other terms are the random fluctuations. We aim at deriving an approximation of these terms, that only depends on 1) the actual value (72), 2) a single random variable $\epsilon_{R}$ for the real part of the fluctuations and 3) a single random variable $\epsilon_{I}$ for the imaginary part of the fluctuations.
As a simple approximation, we separately consider the real and imaginary parts of the fluctuations and then just add them in our final approximate model. The real part is handled as follows. The random variable $\epsilon_{R}$ takes both $\epsilon_{k}$ and $\epsilon_{\ell}$ of (74) into account. Therefore, it represents an error for a ket component and it is drawn with the same statistics as the fluctuations of a ket component of Section V, i.e. it is real-valued and uniformly drawn over the range $[-w/2,w/2]$. In other words, to obtain a simple model, both $\epsilon_{k}$ and $\epsilon_{\ell}$ are here replaced by $\epsilon_{R}$ in (74) (without taking their possible statistical independence into account). Then, to express this real-valued part of the fluctuation model only with respect to $\rho_{k\ell}$ in addition to $\epsilon_{R}$, as explained above, we ignore the phases of $c_{k}$ and $c_{\ell}^{*}$ in the second and third terms of (74) and we replace both of them with $\sqrt{|\rho_{k\ell}|}$, based on (72). The resulting approximate model for the real part of fluctuations, corresponding to the real part of the last three terms of (74), is thus equal to $2\sqrt{|\rho_{k\ell}|}\epsilon_{R}+\epsilon_{R}^{2}$.
As explained above, the imaginary part of fluctuations is handled in the same way and just added to the above real part. As an overall result, the complete approximate noisy model for any density matrix element reads
$\widehat{\rho}_{k\ell}=\rho_{k\ell}+2\sqrt{|\rho_{k\ell}|}\epsilon_{R}+\epsilon_{R}^{2}+i\left(2\sqrt{|\rho_{k\ell}|}\epsilon_{I}+\epsilon_{I}^{2}\right).$ (75)
We again stress that this model contains approximations but meets the required contraints: it allows one to create a “noisy” version $\widehat{\rho}_{k\ell}$ of a given “noiseless” density matrix element $\rho_{k\ell}$ by just drawing a sample of each of the random variables $\epsilon_{R}$ and $\epsilon_{I}$, whose statistics are defined in a way which is qualitatively consistent with those of the random variable used in Section V for deriving a “noisy” version of a ket component. This then mainly allows us to compare the performance of several QPT methods with the same (relevant) fluctuation statistics. This qualitative relevance of our simplified model is thus sufficient for our needs. This is to be contrasted with, e.g., the situation when one would consider a given physical system and one would have to develop a model of it that is as accurate as possible.